### v1

# An overview of meta-analytic methods for economic research

Abstract
Meta-analysis is the use of statistical methods to combine the results of individual studies to estimate the overall effect size for a specific outcome of interest. The direction and magnitude of this estimated effect, along with its confidence interval, provide insights into the phenomenon or relationship being investigated. As an extension of the standard meta-analysis, meta-regression analysis incorporates multiple moderators representing identifiable study characteristics into the meta-analysis model, thereby explaining some of the heterogeneity in true effect sizes across studies. This form of meta-analysis is especially designed to quantitatively synthesize empirical evidence in economics. This study provides an overview of the meta-analytic procedures tailored for economic research. By addressing key challenges, including between-study heterogeneity, publication bias, and effect size dependence, it aims to equip researchers with the tools and insights needed to conduct rigorous and informative meta-analytic studies in economics and related disciplines.
keywords:
Meta-Analysis; Meta-Regression; Heterogeneity; Publication Bias; Effect Size Dependence; Economic Research
journal: arXiv
[…]
footnote
1 Introduction
Meta-analysis, a powerful statistical technique, has become essential for synthesizing research findings across diverse disciplines. By integrating results from multiple studies, it enables researchers to draw comprehensive conclusions that overcome the limitations of individual investigations. The development of meta-analysis is deeply rooted in the history of scientific inquiry, with its origins in early statistical methods of the 17th century and its formalization in the late 20th century.
The foundation for quantitatively synthesizing study results dates back to the 17th century with the work of French mathematician Blaise Pascal, who devised mathematical methods to calculate probabilities in games of chance. These methods were later adopted by astronomers to combine observational data [Cheung & Vijayakumar, 2016].
In the 18th and 19th centuries, mathematicians like Carl Friedrich Gauss and Pierre-Simon Laplace further developed these ideas. George Biddell Airy, the British Astronomer Royal, formalized these methods in his textbook, which influenced later statisticians such as Karl Pearson and Ronald Fisher [O’rourke, 2007]. Karl Pearson In is credited with one of the earliest applications of meta-analytic techniques. In 1904, Pearson conducted a study to examine the relationship between mortality and inoculation with a vaccine for typhoid fever by combining correlation coefficients across multiple studies [Pearson, 1904]. Pearson’s work laid the groundwork for the formalization of meta-analysis, even though his methods lacked the rigor characteristic of modern meta-analyses.
Ronald Fisher, working at the Agricultural Research Station in Rothamsted during the 1920s and 1930s, made significant contributions to the development of meta-analysis. In his 1935 textbook, Fisher illustrated how to effectively analyze multiple agricultural studies, addressing issues such as the variability of fertilizer effects across years and locations [Fisher, 1935]. His colleague, William Cochran, expanded these methods and applied them to both agricultural and medical research, further advancing the techniques of meta-analysis.
The term “meta-analysis” was coined by Gene Glass in 1976 to describe the statistical synthesis of results from multiple individual studies with the goal of integrating their findings [Glass, 1976]. Glass’s work, along with that of his colleagues, drew significant attention to this approach, particularly in the field of psychology. For instance, Mary Smith and Glass published a meta-analysis on the effectiveness of psychotherapy, demonstrating that psychotherapy was effective and that there was little difference in effectiveness across different types of therapies [Smith & Glass, 1977].
The use of meta-analysis in medical research began to take shape in the 1970s. A particularly influential study was the first randomized controlled trial conducted by Peter Elwood, Archie Cochrane, and their colleagues, which aimed to evaluate the effect of aspirin in reducing the recurrence of heart attacks. Although the initial results suggested a potential beneficial effect, they did not achieve statistical significance. As more trials became available, Elwood and Cochrane systematically aggregated and synthesized the findings through meta-analysis, thereby providing a more robust evaluation of aspirin’s efficacy [Elwood et al., 1974].
Richard Peto and his colleagues further advanced the field by providing detailed examples of systematic reviews and meta-analyses, particularly in the context of randomized trials of beta-blockade following heart attacks [Peto et al., 1988]. These contributions were instrumental in promoting the use of meta-analysis in clinical research.
In the decades following its formalization, meta-analysis has become a widely used method across various scientific disciplines, including social sciences, medicine, education, public health, environmental science, psychology, and economics. Notable publications and textbooks, such as Hedges and Olkin’s Statistical Methods for Meta-Analysis (1985), have provided rigorous statistical frameworks for conducting meta-analyses [Hedges & Olkin, 1985].
Overall, meta-analysis has evolved from early statistical methods for combining observations to a sophisticated and widely accepted technique for synthesizing research findings across diverse fields. Its development has been marked by contributions from mathematicians, statisticians, and researchers who recognized the need for systematic approaches to integrating study results. Today, meta-analysis continues to be a vital tool for advancing scientific knowledge and informing evidence-based practice.
Meta-analysis has gained significant importance in economics, offering a structured approach to synthesizing a vast and diverse body of research. Economic studies often yield varying results due to differences in data, methodologies, and contexts. Meta-analysis helps in systematically combining these results, allowing economists to derive more robust and comprehensive conclusions. By integrating findings from multiple studies, meta-analysis enhances statistical power and provides more reliable estimates of effect sizes, which are crucial for understanding economic phenomena [Stanley & Doucouliagos, 2012, Borenstein et al., 2021].
Meta-analysis is a crucial tool in the evaluation of economic policies. Economic policies and interventions are often subject to rigorous analysis, and meta-analysis enables the aggregation of evidence from different studies to assess their overall effectiveness. This approach is particularly valuable in fields such as labour economics, health economics, public economics, environmental economics, resource economics, development economics, international economics, monetary economics, and behavioral economics, where policymakers rely on robust evidence to make informed decisions. For example, meta-analyses have been extensively used to evaluate the effectiveness of labour market policies [Card & Krueger, 1995, Neumark & Wascher, 2006, Card et al., 2010, 2018, Vooren et al., 2019, Campolieti, 2020, Martínez & Martínez, 2021].
Meta-analysis also plays a crucial role in addressing publication bias, a common issue in economic research. Studies with statistically significant or theoretically well-supported findings are more likely to be published, while those with null or insignificant results, or inconsistent with theoretical expectations, are often overlooked. This bias can distort the overall understanding of an economic issue. Meta-analysis techniques, such as funnel plots and statistical tests for publication bias, help in detecting and correcting for these biases, leading to more accurate and credible conclusions [Egger et al., 1997a, Rothstein et al., 2005b].
Furthermore, meta-regression analysis, an extension of meta-analysis, allows economists to explore the heterogeneity in study results by incorporating multiple moderators representing identifiable study characteristics. This technique helps in understanding the factors that contribute to variations in effect sizes across different studies, providing deeper insights into economic relationships. For example, meta-regression can be used to analyze how different labor market conditions or policy environments influence the effectiveness of training programs [Stanley & Doucouliagos, 2012, Borenstein et al., 2021].
In conclusion, meta-analysis has become an indispensable tool in economic research, offering a rigorous and systematic approach to synthesizing a wide range of study results. By addressing issues such as heterogeneity, publication bias, and effect size dependence, meta-analysis provides economists with the tools necessary to conduct comprehensive and informative analyses. This study aims to provide an overview of the meta-analytic procedures tailored for economic research, equipping researchers with the insights needed to conduct robust and impactful meta-analytic studies.
In the following sections, this paper will present a detailed exploration of various meta-analytic methods employed in economic research. Section 2 delves into the methodologies for estimating effect sizes and confidence intervals in meta-analysis, alongside an examination of the different approaches for assessing and quantifying heterogeneity across studies. Section 3 expands on meta-analysis through meta-regression analysis, which incorporates study characteristics (moderators) to address heterogeneity, offering insights into fixed and random effects models, estimation tech-niques, tests for residual heterogeneity, and the influence of moderators on observed effects. Section 4 addresses the challenges of publication bias in meta-analyses, including outcome reporting bias and language bias, and outlines methods such as funnel plots and statistical tests to detect and mitigate these biases. Section 5 introduces the Unrestricted Weighted Least Squares (UWLS) estimator for meta-analysis and meta-regression models, highlighting its advantages over traditional fixed effects models by accounting for residual heterogeneity. Section 6 examines the issue of effect size dependence in meta-analysis, providing a comprehensive analysis of its implications and presenting strategies to effectively manage this challenge. Finally, Section 7 presents concluding remarks.
2 Meta-Analysis
Meta-analysis is a statistical method used to evaluate and combine findings from multiple studies addressing a common research question, with the goal of drawing overall conclusions. The standard meta-analysis process consists of four main steps: (i) selecting an appropriate measure of effect size, (ii) calculating the effect size for each study included in the meta-analysis, (iii) computing the mean effect size as a summary statistic by combining the individual effect sizes, and (iv) performing statistical inference about the true (population) mean effect, including confidence intervals and hypothesis testing, to draw overall conclusions.
The first step in conducting a meta-analysis is to standardize the results from different studies into a common metric, referred to as the “effect size”. This process enables the comparison and aggregation of findings from studies that utilize varying measures of the same construct. For instance, the partial correlation coefficient is commonly used as an effect size measure. It quantifies the strength and direction of the relationship between two quantitative variables while controlling for other variables, and its values range from -1 to +1.111The partial correlation coefficient is widely employed in empirical economic research as a measure of effect size. It is calculated using the following formula: $r_{i}=\frac{t_{i}}{\sqrt{t_{i}^{2}+df_{i}}}$ with standard errors given by: $S_{i}=\sqrt{\frac{1-r_{i}^{2}}{df_{i}}},$ where $t_{i}$ represents the $t$-statistic for the $i$th estimate ($i=1,2,...,k$) reported in the primary studies, and $df_{i}$ denotes the corresponding degrees of freedom [Greene, 2012]. For studies that report $z$-statistics (instead of $t$-statistics) for regression coefficients, the partial correlation coefficient can be derived as $r=\sqrt{z^{2}/n}=z/\sqrt{n}$, where $n$ is the sample size [Rosenthal, 1991]. Unfortunately, the sampling distribution of the correlation coefficient $r$ becomes increasingly skewed as the absolute value of the population effect size ($\rho$) increases [Hedges & Olkin, 1985, Rosenthal, 1991]. To render the variance of $r$ independent of its population counterpart and to normalize its distribution, it is advisable to apply Fisher’s $z$-transformation, as specified by: $z_{i}=\frac{1}{2}\ln\left(\frac{1+r_{i}}{1-r_{i}}\right),$ where $\ln$ denotes the natural logarithm. The corresponding transformation for the population correlation coefficient $\rho_{i}$ is: $\zeta_{i}=\frac{1}{2}\ln\left(\frac{1+\rho_{i}}{1-\rho_{i}}\right).$ In this context, $z_{i}$ represents the Fisher-transformed correlation coefficient, which is approximately normally distributed with a mean of $\zeta_{i}$ and a variance of $\frac{1}{n_{i}-3-p_{i}}$, irrespective of the value of $\zeta_{i}$. Here, $n_{i}$ refers to the sample size associated with the $i$-th effect estimate, and $p_{i}$ is the number of variables conditioned on (see Hedges & Olkin, 1985). For ease of interpretation, the transformation can be reversed, with $z$ converted back to $r$ using the formula: $r=\frac{e^{2z}-1}{e^{2z}+1},$ where $e^{x}=\exp(x)$ denotes the exponential function. However, both Hunter et al. [1996] and Hunter & Schmidt [2004] caution against the indiscriminate use of Fisher’s $z$-transformation. They argue that while the transformation mitigates a negative bias in the untransformed correlation coefficients, it introduces a positive bias in the transformed values. Notably, this bias is generally larger (in absolute magnitude) than the original bias in the untransformed coefficients, particularly when population correlation coefficients exhibit substantial variation across studies. Other examples include the (log) odds ratio, (log) risk ratio, raw mean difference, and standardized mean difference (see Hox et al., 2010).
After obtaining estimates of the individual effect sizes, the subsequent step is to calculate an estimate of the overall (or combined) effect size. All meta-analytic procedures derive the overall effect size by computing a weighted average of the individual effect size estimates, with the primary distinction between methods lying in the specific weighting scheme employed. Meta-analysis typically involves two primary approaches: fixed-effect models and random-effects model [Hedges & Vevea, 1998, Hunter & Schmidt, 2000, Schmidt et al., 2009]. These approaches are outlined in the following discussion.
Consider a meta-analysis involving $k$ independent studies, where study $i$ ($i=1,2,\dots,k$) provides an estimate $y_{i}$ of the corresponding population effect size $\beta_{i}$, and this estimate has an associated within-study sampling variance $\sigma_{i}^{2}$. The model for this situation can be expressed as:
$y_{i}=\beta_{i}+\varepsilon_{i},\quad i=1,2,\dots,k,$ (1)
where the sampling errors $\varepsilon_{i}$ are assumed to be independently distributed as $N(0,\sigma_{i}^{2})$, with $\sigma_{i}^{2}$ typically assumed to be fixed and known. The fixed-effect and random-effects models differ in their treatment of $\beta_{i}$ in this general model. The fixed-effect meta-analysis model is derived under the assumption that $\beta_{i}=\mu_{F}$, a common (or fixed) effect size. Specifically,
$y_{i}=\mu_{F}+\varepsilon_{i},\quad\varepsilon_{i}\sim N(0,\sigma_{i}^{2}),$ (2)
from which it follows that the $y_{i}$’s are independently distributed as $N(\mu_{F},\sigma_{i}^{2})$. In the context of the fixed-effect model, all studies included in the meta-analysis are assumed to share a common true effect size $\mu_{F}$. That is, all studies estimate the same true effect size, and variation in the observed effect sizes arises solely due to random sampling error (within-study variance $\sigma_{i}^{2}$).
In this framework, the goal of the meta-analysis is to estimate the common effect size $\mu_{F}$, accompanied by a confidence interval. A natural estimate of $\mu_{F}$ is the pooled estimate, which can be obtained by computing a weighted average of the individual effect estimates:
$\hat{\mu}_{F}=\frac{\sum_{i=1}^{k}\hat{w}_{i}y_{i}}{\sum_{i=1}^{k}\hat{w}_{i}},\quad\text{with}\quad\widehat{Var}(\hat{\mu}_{F})=\frac{1}{\sum_{i=1}^{k}\hat{w}_{i}},$ (3)
where the “true” weights are the inverse of the within-study sampling variances, $w_{i}=1/\sigma_{i}^{2}$, and $\widehat{Var}(\hat{\mu}_{F})$ is the estimated variance of $\hat{\mu}_{F}$. Thus, the weight assigned to each effect size estimate is inversely proportional to its variance $\sigma_{i}^{2}$. This weighting scheme ensures that effect size estimates with smaller variances (and thus higher precision) receive larger weights in the calculation of the overall estimate, thereby having a greater influence on the resulting combined effect size.
If the within-study sample sizes are sufficiently large (e.g., $n_{i}\geq 30$ for each study, as suggested by Raudenbush & Bryk, 2002), it is reasonable to assume that the effect size estimates are approximately normally distributed with within-study sampling variances, $\sigma_{i}^{2}$, which are estimated in practice by the sample variances $S_{i}^{2}$ from primary studies, though they are conventionally treated as fixed and known in the analysis. Thus, $\hat{w}_{i}=1/S_{i}^{2}$ is used in eq. 3 as an estimate for $w_{i}=1/\sigma_{i}^{2}$. The estimator $\hat{\mu}_{F}$ is equivalent to the weighted least squares (WLS) estimator with weights $\hat{w}_{i}=1/S_{i}^{2}$ (which corresponds to applying ordinary least squares (OLS) to the transformed model $y_{i}/S_{i}=\mu(1/S_{i})+\varepsilon_{i}/S_{i}$). Assuming that the effect size estimates (or $\varepsilon_{i}$’s) are normally distributed, a Wald-type $100(1-\alpha)\%$ confidence interval for the common effect size $\mu_{F}$ can be constructed as $\hat{\mu}_{F}\pm z_{\alpha/2}\sqrt{\widehat{Var}(\hat{\mu}_{F})}$, where $z_{\alpha/2}$ is the upper $\alpha/2$ quantile of the standard normal distribution [Van Houwelingen et al., 2002, Borenstein et al., 2021].
In contrast, under the random effects meta-analysis model, the assumption of a common effect size across all studies is relaxed. Instead, the effect size parameters ($\beta_{1},\beta_{2},\dots,\beta_{k}$) are assumed to be a random sample from a population following a normal distribution $N(\mu_{R},\tau^{2})$. Specifically, the $\beta_{i}$’s are independent and identically distributed random variables, with a common mean $\mu_{R}$ and variance $\tau^{2}$, where $\mu_{R}$ represents the mean of the true effect sizes, and $\tau^{2}$ is the between-study variance, indicating the variability in true effects across studies. Thus, the true effects still share a common mean.
Assuming that $\beta_{i}=\mu_{R}+u_{i}$, the model can be reformulated as:
$y_{i}=\mu_{R}+u_{i}+\varepsilon_{i},\qquad u_{i}\sim N(0,\tau^{2})\;\;\text{and}\;\;\varepsilon_{i}\sim N(0,\sigma_{i}^{2}),$ (4)
where the error terms $u_{i}$ and $\varepsilon_{i}$ are assumed to be uncorrelated. This formulation implies that the $y_{i}$’s follow independent distributions as $N(\mu_{R},\sigma_{i}^{2}+\tau^{2})$. Consequently, the variation in the observed effect sizes is attributed not only to random sampling error (within-study variance, $\sigma_{i}^{2}$) but also to true variation in effect sizes (between-study variance, $\tau^{2}$). The objective of random-effects meta-analysis is, therefore, not to estimate a single common effect across all studies. Instead, it aims to estimate both the mean of the distribution of true effects ($E(\beta_{i})=\mu_{R}$) and the degree of heterogeneity among these effects ($\tau^{2}$).
The random-effects estimator of $\mu_{R}$, like the fixed-effect estimator, is a weighted average of the observed effect sizes (or WLS), expressed as:
$\hat{\mu}_{R}=\frac{\sum_{i=1}^{k}\hat{w}_{i}^{\ast}y_{i}}{\sum_{i=1}^{k}\hat{w}_{i}^{\ast}},\qquad\qquad\text{with}\qquad\widehat{Var}(\hat{\mu}_{R})=\frac{1}{\sum_{i=1}^{k}\hat{w}_{i}^{\ast}},$ (5)
where the weights are the inverse of the variance of the observed effect sizes under the random-effects assumption. Specifically, $w_{i}^{\ast}=1/\sigma_{i}^{2}+\tau^{2}$.
The random-effects meta-analysis model incorporates between-study variance (also referred to as excess variance), $\tau^{2}$, into the weights. This approach reduces to the fixed-effect meta-analysis model when there is no heterogeneity among the true effects (i.e., $\tau^{2}=0$), such that any differences in the observed effect sizes are solely attributed to within-study sampling error. As in the fixed-effect model, the within-study variances $\sigma_{i}^{2}$ are estimated by $S_{i}^{2}$. Although these estimates are used in practice, $\sigma_{i}^{2}$ are typically treated as fixed and known for analytical purposes.
The estimator $\hat{\mu}_{R}$ is equivalent to the WLS estimator, with the weights given by $\hat{w}_{i}^{\ast}=1/(S_{i}^{2}+\hat{\tau}^{2})$.
Several procedures are available for estimating the between-study variance parameter, $\tau^{2}$, which serves as a measure of heterogeneity in the true effects (see Sidik & Jonkman, 2007, Sánchez-Meca & Marín-Martínez, 2008, Veroniki et al., 2016, Langan et al., 2017, Petropoulou & Mavridis, 2017). Among these, the most widely used is the method of moments (MM) estimator proposed by DerSimonian & Laird [1986]. This approach is both computationally simple and non-iterative, and it is calculated as:
$\hat{\tau}^{2}=\max\left\{0,\frac{Q-(k-1)}{\sum_{i=1}^{k}\hat{w}_{i}-\frac{\left(\sum_{i=1}^{k}\hat{w}_{i}\right)^{2}}{\sum_{i=1}^{k}\hat{w}_{i}}}\right\},$ (6)
where $\hat{w}_{i}=1/S_{i}^{2}$ represents the weight for the $i$th effect estimate under a fixed-effect meta-analysis model. The term $Q$, Cochran’s heterogeneity statistic, is defined as:
$Q=\sum_{i=1}^{k}\hat{w}_{i}(y_{i}-\hat{\mu}_{F})^{2}=\sum_{i=1}^{k}\left(\frac{y_{i}-\hat{\mu}_{F}}{S_{i}}\right)^{2},$ (7)
and approximately follows a $\chi^{2}$ distribution with $k-1$ degrees of freedom under the null hypothesis that the true effects are homogeneous ($\tau^{2}=0$) (see Biggerstaff & Tweedie, 1997). The resulting estimate $\hat{\tau}^{2}$ is incorporated into the weighting factor $\hat{w}_{i}^{\ast}=1/S_{i}^{2}+\hat{\tau}^{2}$, which serves as an estimate for the true weight, $w_{i}^{\ast}=1/(\sigma_{i}^{2}+\tau^{2})$, in eq. 5.
In the context of the standard DerSimonian-Laird random-effects model, a $100(1-\alpha)\%$ Wald-type confidence interval for the mean effect size ($\mu_{R}$) is constructed as $\hat{\mu}_{R}\pm z_{\alpha/2}\sqrt{\widehat{Var}(\hat{\mu}_{R})}$, where $z_{\alpha/2}$ represents the upper $\alpha/2$ quantile of the standard normal distribution. However, it has been observed that this confidence interval, based on the normal approximation, is often too narrow (with empirical coverage probabilities lower than the nominal value, particularly when the number of studies is small). This issue arises because the method does not account for the sampling variability (uncertainty) in estimating the heterogeneity parameter ($\hat{\tau}^{2}$). Furthermore, the approach assumes that within-study variances ($\sigma_{i}^{2}$) are fixed and known, an assumption that becomes untenable with small or moderate sample sizes.
To address these limitations, several alternatives to the standard DerSimonian-Laird method have been proposed to incorporate the uncertainty in estimating the between-study variance ($\tau^{2}$) (see Brockwell & Gordon, 2001, 2007, Sidik & Jonkman, 2002, 2006, Ioannidis et al., 2007, Sánchez-Meca & Marín-Martínez, 2008, Cornell et al., 2014, Röver et al., 2015, Bodnar et al., 2017, Guolo & Varin, 2017).
One common approach is to construct confidence intervals for the true mean effect using the $t$-distribution rather than the normal distribution, as $\hat{\mu}_{R}\pm t_{df,\alpha/2}\sqrt{\widehat{Var}(\hat{\mu}_{R})}$, where $t_{df,\alpha/2}$ represents the upper $\alpha/2$ quantile of the $t$-distribution with $df$ degrees of freedom. Determining an effective number of degrees of freedom ($df$) for the distribution of the test statistic $T=\left(\hat{\mu}_{R}-\mu_{R}\right)/\sqrt{\widehat{Var}(\hat{\mu}_{R})}$ is challenging, as it depends on the number of studies included in the meta-analysis ($k$), the extent of between-study heterogeneity, and the magnitude of within-study variances [Higgins et al., 2009]. Consequently, researchers have suggested several degrees of freedom for this test statistic, such as $k-1$ [Follmann & Proschan, 1999], $k-2$ [Raghunathan & Li, 1993, Higgins et al., 2009], and $k-4$ [Berkey et al., 1995].
Alternatively, Hartung & Knapp [2001a, b] introduced a modification to the DerSimonian-Laird method to construct an improved confidence interval for the true mean effect by (i) deriving a refined variance estimator adjusted for small-sample bias and (ii) replacing the normal approximation with a $t$-distribution with $k-1$ degrees of freedom. The test statistic in this approach is given by
$t_{k-1}=\frac{\hat{\mu}_{R}-\mu_{R}}{\sqrt{\widehat{Var}_{S}(\hat{\mu}_{R})}},$ (8)
where $\widehat{Var}_{S}(\hat{\mu}_{R})$ is the refined variance estimator defined as
$\widehat{Var}_{S}(\hat{\mu}_{R})=q\cdot\widehat{Var}(\hat{\mu}_{R}),\quad\text{with}\quad q=\left\{\frac{1}{k-1}\sum_{i=1}^{k}\hat{w}_{i}^{\ast}(y_{i}-\hat{\mu}_{R})^{2}\right\}.$ (9)
A $100(1-\alpha)\%$ confidence interval for the true mean effect $\mu_{R}$ would then be calculated as $\hat{\mu}_{R}\pm t_{k-1,\alpha/2}\sqrt{\widehat{Var}_{S}(\hat{\mu}_{R})}$, where $t_{k-1,\alpha/2}$ is the upper $\alpha/2$ quantile of the $t$-distribution with $k-1$ degrees of freedom. This procedure was independently proposed by Sidik & Jonkman [2002]. The so-called Hartung-Knapp-Sidik-Jonkman method typically produces confidence intervals that are wider than those based on the standard normal distribution. However, if $q$ is sufficiently small, the modified confidence interval may paradoxically become shorter than the standard one. To address this limitation, Röver et al. [2015] proposed a simple modification: replacing $q$ in eq. 9 with $q^{\ast}=\max\{1,q\}$. It is worth noting that the estimator of $\tau^{2}$ used in this method is structurally identical to that employed in the DerSimonian-Laird approach. Nevertheless, the proposed confidence intervals are relatively robust to variations in the procedure used to estimate $\tau^{2}$ [Sidik & Jonkman, 2003, 2006].
Additionally, Sidik & Jonkman [2006] proposed an alternative, approximately unbiased estimator for the variance of the overall effect estimate ($\hat{\mu}_{R}$) in random effects meta-analysis. This estimator is robust to errors in the estimated marginal variances ($S_{i}^{2}$ and $\hat{\tau}^{2}$) that underlie the random effects weights. The authors advocate for using the robust variance estimator or the weighted sample variance estimator instead of the conventional variance estimator when making inferences about the overall effect size in the random effects meta-analysis framework.
Another approach is the method proposed by Kenward & Roger [1997], which adjusts the variance of the overall effect estimate using the expected information and modifies the degrees of freedom of the $t$-distribution accordingly. This method provides an alternative for constructing confidence intervals and testing hypotheses for the overall effect size (see Partlett & Riley, 2017).
Several alternative estimators for $\tau^{2}$ have also been developed to incorporate the uncertainty associated with its estimation in the standard DerSimonian-Laird approach (see Veroniki et al., 2016). For instance, the likelihood ratio (LR) method described by Hardy & Thompson [1996] employs a profile likelihood function to simultaneously estimate $\tau^{2}$ and $\mu_{R}$, as well as to construct likelihood-based confidence intervals.
Partlett & Riley [2017] conducted a simulation study to compare the coverage probabilities of confidence intervals generated by these methods, specifically under restricted maximum likelihood (REML) estimation of $\tau^{2}$ instead of the standard DerSimonian-Laird estimation. Their findings suggest that the Hartung-Knapp-Sidik-Jonkman method outperforms other approaches when between-study heterogeneity is large and/or within-study sample sizes are similar. However, ven this method is problematic when heterogeneity is low, and within-study sample sizes vary significantly.
The primary objective of the meta-analysis is to estimate the mean effect size and its confidence interval. The confidence interval provides an indication of the precision of the estimate of the mean effect size (based on the standard error of the mean), as it shows the range within which the true mean effect is likely to lie [Lipsey & Wilson, 2001]. Ideally, this interval should be as narrow as possible (the narrower the interval, the more precise the estimate). However, in a random effects meta-analysis, such intervals do not convey information about the dispersion of true effect sizes, which is also of significant interest. The prediction interval is used to address this, as it indicates the range within which the true effect size in a new (or future) study, similar to those included in the meta-analysis, is likely to lie. An approximate $100(1-\alpha)\%$ prediction interval for the new effect can be formed using $\hat{\mu}_{R}\pm t_{k-2,\alpha/2}\sqrt{\widehat{Var}(\hat{\mu}_{R})+\hat{\tau}^{2}}$, where $t_{k-2,\alpha/2}$ is the upper $\alpha/2$ quantile of the $t$-distribution with $k-2$ degrees of freedom [Higgins et al., 2009]. Partlett & Riley [2017] conducted a simulation study to compare the coverage of several versions of this so-called Higgins-Thompson-Spiegelhalter prediction interval, constructed using different estimators for the variance of the mean effect estimate (Hartung-Knapp-Sidik-Jonkman, Sidik-Jonkman, and Kenward-Roger variance estimators), following REML estimation of $\tau^{2}$. They conclude that all procedures perform reasonably well when $k\geq 5$ and the degree of heterogeneity is sufficiently large. It is evident that the uncertainty in the estimate of $\tau^{2}$ affects both terms in the estimated variance of the new effect drawn from the true random effects distribution, $\left[\widehat{Var}(\hat{\mu}_{R})+\hat{\tau}^{2}\right]$. Nagashima et al. [2019] propose a parametric bootstrap approach to construct prediction intervals that account for this uncertainty and are valid even when the number of studies included in the meta-analysis ($k$) is small. Using Monte Carlo simulations, they compare the coverage probabilities of the proposed prediction interval with the Higgins-Thompson-Spiegelhalter prediction intervals based on the standard, Hartung-Knapp-Sidik-Jonkman, and Sidik-Jonkman variance estimators for $\hat{\mu}_{R}$, following REML estimation of $\tau^{2}$. Their results indicate that the coverage performance of the Higgins-Thompson-Spiegelhalter prediction intervals may be poor, particularly when the relative degree of heterogeneity is low or moderate. In contrast, the coverage probability of the proposed prediction interval remains nearly constant and close to the nominal level, regardless of the amount of heterogeneity.
Assessment of statistical heterogeneity is another important aspect of meta-analysis, which may be conducted by testing for the presence of heterogeneity, measuring the extent of heterogeneity, quantifying its impact, or using a combination of these methods [Higgins, 2008]. As discussed by Thompson [1994], the impact of heterogeneity on the overall conclusions of the meta-analysis needs to be considered carefully. Tests for heterogeneity are commonly used to decide which statistical model is appropriate for a particular meta-analysis. If there is evidence of heterogeneity in the true effect sizes, then a random effects model should be used, whereas a fixed effect model is more appropriate when there is no significant heterogeneity across studies. [^2] Hedges & Vevea [1998] point out that the fixed and random effects procedures are designed for different inference goals, implying that the choice of a statistical procedure should be based on the nature of the inference the analyst wishes to make about the effect size parameters. They distinguish between conditional and unconditional inferences. In the former case, the analyst wishes to make statistical inferences about the effect size parameters in the set of studies being analyzed (or a set of studies identical to those included in the analysis except for uncertainty arising from the sampling of participants into the studies) and say nothing about other studies. In this case, the fixed effects model is chosen to treat the effect size parameters as fixed unknown constants (fixed effects) to be estimated. In contrast, the random effects model is appropriate when the analyst wishes to make an explicit generalization beyond the collection of studies being analyzed (referred to as unconditional inference). The random effects procedure treats each observed study’s effect size as a random draw from an underlying population or distribution of effect size parameters (or a superpopulation), and makes inferences about the mean and variance of this distribution. A simple way to assess heterogeneity in the true effect sizes is through the “forest plot.” The forest plot is a graphical tool used to represent the point estimates of the effect sizes (both individual and overall) and the corresponding confidence intervals in a simple visual display. Less overlap between confidence intervals is associated with greater heterogeneity between studies [Lewis & Clarke, 2001, Panesar et al., 2010, Hanji, 2017]. In addition, Cochran’s $Q$ statistic ($\sim\chi_{k-1}^{2}$), as described in eq. 7, is commonly used in meta-analyses to test the null hypothesis of homogeneity — $\text{H}_{0}:\tau^{2}=0$ or $\beta_{i}=\mu_{F}$ for all $i$ — against the alternative of heterogeneity. If $Q$ is greater than the upper $\alpha$ quantile of the $\chi_{k-1}^{2}$ distribution, then the null hypothesis would be rejected at the significance level $\alpha$. If $\text{H}_{0}$ is rejected, the meta-analyst may conclude that the true effects are heterogeneous and a random effects model is more appropriate for the given meta-analysis. Furthermore, the confidence interval provides a useful tool for testing a hypothesis about $\tau^{2}$. If the confidence interval for $\tau^{2}$ does not contain 0, the null hypothesis that $\tau^{2}=0$ is rejected. The confidence interval provides a measure of the degree of uncertainty associated with the point estimate of the amount of heterogeneity. In other words, it is a measure of precision (based on the standard error), which gives an indication of how precisely the heterogeneity parameter, $\tau^{2}$, is estimated. There are a number of different approaches to constructing confidence intervals for $\tau^{2}$. These include Wald-type approaches, likelihood-based approaches, approaches using forms of generalized Cochran’s $Q$ statistics, bootstrap approaches, Bayesian approaches, and others (see Knapp et al., 2006, Viechtbauer, 2007a, Jackson, 2013, Jackson et al., 2015, Jackson & Bowden, 2016, Veroniki et al., 2016).
Having tested for and found evidence of heterogeneity across studies, the logical next step in the meta-analysis is to quantify this heterogeneity. It is important to quantify between-study heterogeneity because the degree of heterogeneity determines the extent to which the findings of the meta-analysis can be generalized, thus influencing the overall conclusions of the meta-analysis [Higgins, 2008]. The $p$-value of the $Q$-test is often quoted as an indication of the degree of heterogeneity. A low $p$-value provides evidence of significant heterogeneity in the effect estimates (beyond that expected due to sampling error alone). However, it has been shown that the statistical power of the $Q$-test for detecting heterogeneity depends on the number and size of studies included in the meta-analysis, as well as the between-study variance parameter, $\tau^{2}$. In particular, this test has poor power when the number of studies or within-study sample sizes is small, and excessive power when there are too many studies, especially with large sample sizes [Sackett et al., 1986, Spector & Levine, 1987, Alexander et al., 1989, Sánchez-Meca & Marín-Martínez, 1997, Hardy & Thompson, 1998, Viechtbauer, 2007b]. Another approach for quantifying heterogeneity is estimating the between-study variance, $\tau^{2}$, as done in a random effects meta-analysis. The problem with this approach is that it depends mainly on the particular effect size metric used in the meta-analysis. Therefore, it is not possible to make comparisons of the degree of heterogeneity across meta-analyses that have used different measures of effect size [Higgins & Thompson, 2002, Huedo-Medina et al., 2006].
Higgins & Thompson [2002] propose a simple measure to quantify between-study heterogeneity in a meta-analysis as follows: [^3] In addition to $I^{2}$, Higgins & Thompson [2002] develop two further measures of the impact of heterogeneity: $H=\sqrt{Q/(k-1)}$ and $R=\sqrt{\left\{\frac{\sum_{i=1}^{k}\hat{w}_{i}}{\sum_{i=1}^{k}\hat{w}_{i}^{\ast}}\right\}}=\sqrt{\left\{\frac{\sum_{i=1}^{k}\hat{w}_{i}}{\sum_{i=1}^{k}(\hat{w}_{i}^{-1}+\hat{\tau}^{2})^{-1}}\right\}}$. They conclude that $H$ and $I^{2}$ are particularly useful measures of the impact of heterogeneity. However, the most popular measure is $I^{2}$, which is a transformation of $H$, defined as $I^{2}=\frac{H^{2}-1}{H^{2}}$.
$I^{2}=\frac{\hat{\tau}^{2}}{S^{2}+\hat{\tau}^{2}},$ (10)
(as an estimate of $\tau^{2}/(\sigma^{2}+\tau^{2})$), where $\hat{\tau}^{2}$ is an estimate of the between-study variance, $\tau^{2}$, which can be obtained by the method of moments as shown in eq. 6, and $S^{2}$ provides an estimate of the so-called “typical” within-study sampling variance, $\sigma^{2}$, which can be derived from the individual within-study variances as follows:
$S^{2}=\frac{(k-1)\sum_{i=1}^{k}\hat{w}_{i}}{\left(\sum_{i=1}^{k}\hat{w}_{i}\right)^{2}-\sum_{i=1}^{k}\hat{w}_{i}^{2}},$ (11)
where $\hat{w}_{i}=1/S_{i}^{2}$ and $k$ is the number of effect estimates. By substituting eq. 6 and eq. 11 into eq. 10, the $I^{2}$ statistic can be rewritten as
$I^{2}=\max\left\{0,\frac{Q-(k-1)}{Q}\right\}.$ (12)
As pointed out by Higgins [2008], $I^{2}$ is not a point estimate of “between-study variance” or a measure of the degree of “between-study heterogeneity”. It represents the proportion of the total variation in observed effect sizes that is due to heterogeneity in the true effect sizes rather than sampling error. That is, the ratio of the between-study variance to the variance of the observed effects (i.e., the sum of the between-study variance and the within-study sampling variance). Borenstein et al. [2017] provide many of the practical details.
$I^{2}$ can also be viewed as a measure of “inconsistency” in the studies’ results in a meta-analysis, since it reflects the degree of overlap of confidence intervals across effect size estimates [Higgins et al., 2003, Higgins, 2008, Borenstein et al., 2021]. Thus, this measure quantifies the impact rather than the extent of between-study heterogeneity in a meta-analysis, ranging from 0% to 100%, with higher values indicating greater impact of true heterogeneity on the conclusions of the meta-analysis. A useful guide for the interpretation of $I^{2}$ is as follows: 0-40% may not be important, 30-60% may represent moderate heterogeneity, 50-90% may represent substantial heterogeneity, and 75-100% represents considerable heterogeneity [Deeks et al., 2008]. In addition, Higgins et al. [2003] suggest that $I^{2}$ values on the order of 25%, 50%, and 75% may be regarded as low, moderate, and high degree of heterogeneity, respectively. A confidence interval for $\tau^{2}/(\sigma^{2}+\tau^{2})$ can be obtained by substituting the endpoints of a confidence interval for $\tau^{2}$ into eq. 10, whereby statistical homogeneity is rejected if the confidence interval does not contain zero. For practical application, Higgins & Thompson [2002] also suggest a simple computation of an interval only based on $Q$ and $k$.
In general, the main advantages of $I^{2}$ are that it (i) is easy to calculate and has an intuitive interpretation, (ii) can be readily calculated for the majority of published meta-analyses, (iii) is characterized by an uncertainty interval, (iv) is inherently independent of the number of studies included in the meta-analysis, (v) can be directly compared across meta-analyses with different types of outcome data and different measures of effect size, and (vi) has a wide range of applications, such as the investigation of the nature and causes of heterogeneity [Higgins & Thompson, 2002, Higgins et al., 2003].
3 Meta-Regression Analysis
In general, as discussed above, the total variation in the observed effect size can be attributed to within-study variance (sampling error variance) and between-study variance (heterogeneity in the true effects). An important aspect of a meta-analysis is identifying potential sources of heterogeneity between studies. Meta-regression has become a widely used tool for addressing this issue. In particular, the fixed and random effects meta-regression models have been developed to investigate whether identifiable study characteristics (moderators) may explain at least part of the heterogeneity among the true effects (see Thompson & Sharp, 1999, Higgins & Thompson, 2004, Harbord & Higgins, 2008, Viechtbauer et al., 2015).
The choice between these two models depends on whether or not all the heterogeneity beyond chance (sampling error) can be systematically accounted for by the moderator variables.
The fixed effects meta-regression model extends the fixed effect meta-analysis model by replacing $\mu_{F}$ with a linear predictor, $\bm{x}_{i}^{\prime}\bm{\beta}_{F}$, in eq. 2:
$y_{i}=\bm{x}_{i}^{\prime}\bm{\beta}_{F}+\varepsilon_{i},\qquad\varepsilon_{i}\sim N(0,\sigma_{i}^{2}),$ (13)
where $\bm{x}_{i}$ is a $(p+1)\times 1$ vector of $p$ explanatory variables (covariates) plus a $1$ in the first position corresponding to the model intercept, and $\bm{\beta}_{F}$ is a $(p+1)\times 1$ vector of unknown constants to be estimated. This formulation implies that the $y_{i}$’s are distributed as $N(\bm{x}_{i}^{\prime}\bm{\beta}_{F},\sigma_{i}^{2})$. To correct for heteroscedasticity, the parameters $\bm{\beta}_{F}=(\beta_{0},\beta_{1},...,\beta_{p})^{\prime}$ can be estimated easily by weighted least squares (WLS) as follows:
$\bm{\hat{\beta}}_{F}=\left[\sum_{i=1}^{k}\hat{w}_{i}\bm{x}_{i}\bm{x}_{i}^{\prime}\right]^{-1}\left[\sum_{i=1}^{k}\hat{w}_{i}\bm{x}_{i}y_{i}\right]=(\bm{X}^{\prime}\bm{\hat{W}}\bm{X})^{-1}(\bm{X}^{\prime}\bm{\hat{W}}\bm{y}),$ (14)
where $\bm{X}$ is a $k\times(p+1)$ matrix of observations with the $i$-th row equal to $\bm{x}_{i}^{\prime}=(1,x_{1i},...,x_{pi})$, $\bm{y}=(y_{1},...,y_{k})^{\prime}$ is a $k\times 1$ vector of the observed effect sizes, and $\bm{\hat{W}}$ is a $k\times k$ diagonal weighting matrix whose $i$-th diagonal element is $\hat{w}_{i}=\frac{1}{S_{i}^{2}}$.
In general, the variance-covariance matrix of this WLS estimator is given by
$\widehat{Var}(\bm{\hat{\beta}}_{F})=\left(\bm{X}^{\prime}\hat{\bm{W}}\bm{X}\right)^{-1}\left(\bm{X}^{\prime}\hat{\bm{W}}\bm{\Sigma}\hat{\bm{W}}\bm{X}\right)\left(\bm{X}^{\prime}\hat{\bm{W}}\bm{X}\right)^{-1},$ (15)
where $\bm{\Sigma}$ is the variance-covariance matrix of the effect size estimates. The diagonal elements of $\bm{\Sigma}$ are the variances associated with the effect sizes, and the off-diagonal elements are the covariances. Assuming that the effect size estimates are statistically independent (i.e., the off-diagonal elements are all equal to zero), and that inverse variance weights are used, then
$\widehat{Var}(\bm{\hat{\beta}}_{F})=(\bm{X}^{\prime}\bm{\hat{W}}\bm{X})^{-1},$ (16)
where the $j$-th diagonal element represents the variance of the estimator $\hat{\beta}_{j-1}$.
The fixed effects meta-regression model assumes that between-study heterogeneity can be fully explained by the covariates $X_{1},...,X_{p}$. This model is generally not recommended because it leads to unacceptably inflated Type I error rates when there is unexplained, or residual, heterogeneity among the true effects, which is often the case in practical applications [Higgins & Thompson, 2004].
The random effects meta-regression model is obtained by replacing $\mu_{R}$ with $\bm{x}_{i}^{\prime}\bm{\beta}_{R}$ in eq. 4:
$y_{i}=\bm{x}_{i}^{\prime}\bm{\beta}_{R}+u_{i}+\varepsilon_{i},\qquad u_{i}\sim N(0,\tau_{res}^{2})\quad\text{and}\quad\varepsilon_{i}\sim N(0,\sigma_{i}^{2}),$ (17)
which implies that the $y_{i}$’s are distributed as $N(\bm{x}_{i}^{\prime}\bm{\beta}_{R},\sigma_{i}^{2}+\tau_{res}^{2})$. This model is referred to as the mixed effects meta-analysis model (the meta-regression coefficients $\beta_{0},\beta_{1},...,\beta_{p}$ are fixed effects, and the $u_{i}$’s are mutually independent random effects). It assumes that the variation in the observed effect size beyond that expected by sampling error (i.e., the variation in the true effect size) has both systematic and random components. The systematic variation is explained by the moderator variables included in the meta-regression model, while the random variation, known as the residual between-study variance, remains unexplained. In other words, the mixed effects meta-analysis model assumes that between-study heterogeneity ($\tau^{2}$) is partitioned into two components: unexplained or residual heterogeneity ($\tau_{res}^{2}$) and explained heterogeneity ($\tau^{2}-\tau_{res}^{2}$). [^4] Note that the variance $\tau_{res}^{2}$ is the conditional variance of the true effect sizes given the covariates. Therefore, all models discussed earlier are restricted versions of the mixed effects meta-analysis model.
For the mixed effects meta-analysis model, it is necessary to estimate both the regression coefficients and the amount of residual heterogeneity in the true effect sizes. Similar to the fixed effects meta-regression model, the estimates of $\bm{\beta}_{R}$ and the corresponding variance-covariance matrix can be obtained using the weighted least squares estimator. The only difference is that, for the random meta-regression model, the weights are $\hat{w}_{i}^{\ast}$ instead of $\hat{w}_{i}$ as in eqs. 14 and 16. Thus,
$\bm{\hat{\beta}}_{R}=\left[\sum_{i=1}^{k}\hat{w}_{i}^{\ast}\bm{x}_{i}\bm{x}_{i}^{\prime}\right]^{-1}\left[\sum_{i=1}^{k}\hat{w}_{i}^{\ast}\bm{x}_{i}y_{i}\right]=(\bm{X}^{\prime}\bm{\hat{W}}^{\ast}\bm{X})^{-1}(\bm{X}^{\prime}\bm{\hat{W}}^{\ast}\bm{y}),$ (18)
and
$\widehat{Var}(\bm{\hat{\beta}}_{R})=(\bm{X}^{\prime}\bm{\hat{W}}^{\ast}\bm{X})^{-1},$ (19)
where the weighting matrix is now $\bm{\hat{W}}^{\ast}=\text{diag}[\hat{w}_{1}^{\ast},\dots,\hat{w}_{k}^{\ast}]$ and $\hat{w}_{i}^{\ast}=1/(S_{i}^{2}+\hat{\tau}_{res}^{2})$, with $\hat{\tau}_{res}^{2}$ being an estimate of the amount of between-study variance $\tau_{res}^{2}$, that is, the degree of heterogeneity in the true effects not accounted for by the moderators included in the model. Many algorithms have been developed for the estimation of this parameter [Thompson & Sharp, 1999, Raudenbush, 2009, Panityakul et al., 2013, Chung et al., 2013, López-López et al., 2014, Viechtbauer et al., 2015, van Aert & Jackson, 2018]. The most common ones include the method of moments (MM), maximum likelihood (ML), restricted maximum likelihood (REML), and empirical Bayes (EB) estimators.
In the context of a meta-regression, meta-analysts can use a generalized version of Cochran’s $Q$-test, as described in eq. 7, to choose between the fixed and random/mixed models. The test is based on the residual heterogeneity statistic
$Q_{res}=\sum_{i=1}^{k}\hat{w}_{i}(y_{i}-\bm{x}_{i}\bm{\hat{\beta}}_{F})^{2}=\sum_{i=1}^{k}\left(\frac{y_{i}-\bm{x}_{i}\bm{\hat{\beta}}_{F}}{S_{i}}\right)^{2},$ (20)
which is distributed as $\chi^{2}$ with $k-p-1$ degrees of freedom under the null hypothesis of no residual/unexplained heterogeneity ($\tau_{res}^{2}=0$). 555 The method of moments (MM) estimator for $\tau^{2}$ in the context of a meta-regression is $\hat{\tau}_{res}^{2}=\frac{Q_{res}-(k-p-1)}{tr(\bm{M})},$ where $tr(\bm{M})$ is the trace of the matrix $\bm{M}=\bm{\hat{W}}-\bm{\hat{W}}\bm{X}(\bm{X}^{\prime}\bm{\hat{W}}\bm{X})^{-1}\bm{X}^{\prime}\bm{\hat{W}},$ (see Raudenbush, 2009, López-López et al., 2017).
If $Q_{res}$ is statistically significant, then some of the between-study heterogeneity remains even after accounting for study characteristics as moderators. In other words, even after accounting for moderators, the true effect sizes remain heterogeneous [Lipsey & Wilson, 2001, Borenstein et al., 2021]. In this case, the mixed effects meta-analysis model would be preferred. However, as previously mentioned, it may be more useful to quantify the heterogeneity in the true effects rather than rely on an overall test to detect its presence [Viechtbauer, 2007a, Harbord & Higgins, 2008].
In addition to the amount of residual heterogeneity, $\hat{\tau}_{res}^{2}$, a measure quantifying the impact of residual heterogeneity on the results can be defined as follows:
$I_{res}^{2}=\frac{\hat{\tau}_{res}^{2}}{S^{2}+\hat{\tau}_{res}^{2}}=\max\left\{0,\frac{Q_{res}-(k-p-1)}{Q_{res}}\right\}.$ (21)
This statistic represents the proportion of variation in the observed effect sizes, beyond what is explained by the covariates, that is attributable to residual (unexplained) heterogeneity rather than sampling error. It ranges from 0% to 100%, with higher values indicating a greater influence of unexplained heterogeneity on the conclusions of the meta-regression analysis.
As in random effects meta-analysis, it is essential to construct confidence intervals for the between-study variance within the context of a meta-regression model (i.e., the residual heterogeneity parameter, $\tau_{res}^{2}$) as an alternative method of inference for the heterogeneity parameter. Jackson et al. [2014] present two exact methods for constructing confidence intervals for $\tau_{res}^{2}$: confidence intervals based on a form of generalized Cochran $Q$ statistics and the $Q$ profile method. They also advocate for the use of informative prior distributions for this parameter in Bayesian meta-regression. All three methods extend the approaches proposed for random effects meta-analysis. Furthermore, Jackson et al. [2015] propose a straightforward method for constructing approximate confidence intervals for the between-study variance component (heterogeneity) that can be applied to moment-based estimators under both random effects meta-analysis and meta-regression models.
An additional index for assessing the impact of moderators, pseudo $R^{2}$, is defined as follows:
$R^{2}=\frac{\hat{\tau}^{2}-\hat{\tau}_{res}^{2}}{\hat{\tau}^{2}}=1-\frac{\hat{\tau}_{res}^{2}}{\hat{\tau}^{2}},$ (22)
where $\hat{\tau}^{2}$ represents the estimate of total between-study heterogeneity derived from the random effects meta-analysis model (refer to eq. 4), and $\hat{\tau}_{res}^{2}$ represents the estimate of residual (unexplained) heterogeneity derived from the mixed effects meta-analysis model (refer to eq. 17). This index quantifies the proportion of the total between-study heterogeneity explained by the moderators included in the mixed effects meta-analysis model. It ranges from 0 to 1 (or 0% to 100%), with higher values indicating that the moderators explain a larger share of the heterogeneity in the true effects [Pastor & Lazowski, 2018, Borenstein et al., 2021].
Several procedures have been proposed in the literature for testing the statistical significance of individual moderator effects and constructing corresponding confidence intervals. These include standard (Wald-type) tests based on the $t$ and $z$ distributions, the likelihood ratio ($LR$) test, permutation tests, resampling tests, the Knapp-Hartung method (an extension of the Hartung-Knapp-Sidik-Jonkman method), and the robust (Huber-White-type) method (see Knapp & Hartung, 2003, Higgins & Thompson, 2004, Raudenbush, 2009, Huizenga et al., 2011, Viechtbauer et al., 2015).
Additionally, determining whether the regression model as a whole is statistically significant is an important question. This can be addressed by testing the joint hypothesis that all coefficients, excluding the constant term, are zero ($\beta_{1}=\dots=\beta_{p}=0$). A common procedure for this is the omnibus (Wald-type) test. The test statistic is computed as $Q_{m}=\bm{\hat{\beta}}^{\prime}_{p}\hat{V}^{-1}\bm{\hat{\beta}}_{p}$, where $\bm{\hat{\beta}}_{p}=(\hat{\beta}_{1},\dots,\hat{\beta}_{p})^{\prime}$ is a $p\times 1$ vector of estimated coefficients, and $\hat{V}$ is the estimated variance-covariance matrix of $\bm{\hat{\beta}}_{p}$. Under the null hypothesis that all $p$ slope coefficients in the meta-regression model are zero, the $Q_{m}$ statistic follows a $\chi^{2}$ distribution with $p$ degrees of freedom, which is equal to the difference between the degrees of freedom of $Q$ and $Q_{res}$, expressed as $p=(k-1)-(k-p-1)$. Rejecting the null hypothesis indicates that at least one of the individual regression coefficients is significantly different from zero, rendering the fitted model significant [Lipsey & Wilson, 2001, Borenstein et al., 2021].
An alternative approach to test the joint hypothesis that all coefficients for the moderator variables are zero is the $LR$ or chi-square difference test. The $LR$ test statistic is $LR=-2[\ln L_{1}-\ln L_{0}]$, where $\ln L_{1}$ and $\ln L_{0}$ are the log-likelihood values for the models with and without the imposed restrictions, respectively. The test statistic follows a $\chi^{2}$ distribution with $p$ degrees of freedom, corresponding to the number of restrictions imposed [Greene, 2012]. However, it is important to note that the $LR$ test is invalid when the restricted maximum likelihood (REML) procedure is used for model estimation. This limitation arises because REML only permits comparisons of models with identical fixed parts/regression coefficients [Hox et al., 2010].
The omnibus and $LR$ tests are analogous to the $F$ test, which examines whether all slope coefficients in a regression model are zero. These tests can also be applied to assess whether one or a subset of coefficients for the moderator variables in a multiple meta-regression model are zero.
4 Publication Bias
Publication selection bias occurs when the selection of studies for publication is influenced by the statistical significance and/or direction of the results, such that studies reporting statistically significant results and/or those consistent with theoretical expectations are more likely to be published than studies with insignificant or unexpected results [Begg & Berlin, 1989, Rothstein et al., 2005b, Doucouliagos, 2005, Scherer et al., 2007, Sutton, 2009, Jennions et al., 2013, Stanley et al., 2017, Borenstein et al., 2021, Page et al., 2022]. Unfortunately, publication bias is prevalent across most research areas [Stanley & Doucouliagos, 2013]. This phenomenon is commonly referred to as the “file drawer” problem [Rosenthal, 1979], with file drawers metaphorically representing the location of the missing/unpublished studies. Because the missing studies are a non-random subset of all relevant studies (i.e., they systematically differ from the published ones), the studies included in a meta-analysis form a biased, unrepresentative sample of the entire body of research (both published and unpublished). In other words, publication bias leads to the selective and biased inclusion of studies in meta-analyses. This issue presents a significant threat to the validity of meta-analytic reviews and other synthesis methods [Rothstein, 2008, Sutton, 2009, Borenstein et al., 2021]. Moreover, many authors may refrain from submitting studies with statistically insignificant and/or inconsistent (i.e., wrong-signed) effects for publication, anticipating rejection by journals [Egger & Smith, 1998]. This issue, known as “outcome reporting bias”, has garnered considerable attention in recent years [Hutton & Williamson, 2000, Hahn et al., 2002, Chan et al., 2004a, b, Williamson & Gamble, 2005, Dwan et al., 2008, Smyth et al., 2011, Pigott et al., 2013, Copas et al., 2013, Jennions et al., 2013, Stanley et al., 2017, Andrews & Kasy, 2019, Bom & Rachinger, 2019, Marks-Anglin & Chen, 2020, Page et al., 2022, van Aert & Wicherts, 2024]. Selective reporting can take several forms, including the selective reporting of certain outcomes from a given study based on the nature and direction of the results, as well as the analytical manipulation of unfavorable results to make them statistically significant or theoretically interesting. This type of bias also contributes to publication bias.
In addition to publication bias, other factors may contribute to bias in the inclusion of individual studies in a meta-analysis. Among published studies, those with significant effects in the expected direction are more likely to be published in English and therefore more likely to be searched (English language bias; Egger et al., 1997b, Egger & Smith, 1998, Jüni et al., 2002), more likely to be published more than once (duplication bias; Easterbrook et al., 1991, Tramèr et al., 1997), more likely to be cited by other authors and therefore easier to find (citation bias; Gøtzsche, 1987, Kjaergard & Gluud, 2002, Ravnskov, 1992), more likely to be published quickly (time-lag bias; Stern & Simes, 1997, Ioannidis, 1998, Jennions & Møller, 2002), and more likely to be published in journals indexed in major electronic databases, particularly in the case of developing countries (database bias; Egger & Smith, 1998, Kjaergard & Gluud, 2002). In addition, published studies that are easily accessible to researchers, available free or at low cost, or limited to one’s own discipline are more likely to be included in a meta-analysis, leading to availability bias, cost bias, and familiarity bias, respectively [Rothstein et al., 2005a, Wade et al., 2006]. There are two strategies to address these and other related biases: conducting comprehensive searches for relevant studies and the prospective registration of research [Song et al., 2002]. These biases fall under the broader umbrella of publication selection bias [Borenstein et al., 2021]. [^6] The combination of publication bias and related selection biases is often referred to as “dissemination bias” [Song et al., 2002].
Several graphical and statistical methods have been proposed for detecting and correcting for publication bias in meta-analysis. The funnel plot is the most commonly used approach for identifying the presence of publication bias and for providing a visual summary of a meta-dataset [Song et al., 2000, Sutton et al., 2000, Sterne et al., 2001, Stanley, 2005, Sterne et al., 2005, Sutton, 2009]. It is a scatter plot of the effect estimates from individual studies on the horizontal axis against a measure of sample size on the vertical axis [Light & Pillemer, 1984, Begg & Berlin, 1988, Light et al., 1994]. Standard error, variance, and their inverses are valid choices for the vertical axis in a funnel plot [Sterne & Egger, 2001], among which the inverse of the standard error (precision) is the most commonly used. Accordingly, the precision of the effect estimate increases as the sample size grows larger (i.e., larger samples provide more precise estimates of effect sizes). In this context, studies with large samples and, therefore, greater precision (smaller standard errors) appear at the top of the plot and tend to cluster near the mean effect size under the fixed-effect meta-analysis model (represented as a vertical line). In contrast, smaller studies with lower precision appear at the bottom of the plot and are more widely distributed because they are subject to greater sampling variation in effect estimates (i.e., the dispersion of effect estimates is inversely correlated with their precision). When there is no publication bias, the resulting plot should be approximately funnel-shaped, hence its name, with its major axis aligned with the mean effect size and the studies symmetrically distributed around this value. However, in the presence of publication bias, the studies are expected to exhibit symmetry at the top, a few missing data points in the middle, and more missing data near the bottom [Doucouliagos et al., 2005, Borenstein et al., 2021]. Under publication selection, researchers conducting smaller-sample studies may need to search more extensively across model specifications, proxies/indices, econometric procedures, and datasets to find estimates large enough to offset their larger standard errors, ensuring that the test statistic reaches a statistically significant level. By contrast, researchers with larger samples and smaller standard errors need not search as rigorously to achieve the desired significance level, as smaller estimates suffice. Consequently, in the presence of publication bias, the reported effect estimates are positively related to their standard errors [Stanley, 2004, Doucouliagos & Stanley, 2009, Doucouliagos et al., 2012]. Furthermore, if the expected sign of the effect is positive, small studies providing small (insignificant) and negative estimates with large standard errors may be missing, resulting in a plot skewed to the right. Conversely, if the expected sign of the effect is negative, small studies providing small (insignificant) and positive estimates with large standard errors may be missing, causing the plot to be skewed to the left.
Doucouliagos [2005], Doucouliagos & Stanley [2013], and Doucouliagos et al. [2018] argue that publication bias is less likely to occur in a contested research field. Specifically, in cases where there is theoretical support for both significant positive and negative effects, researchers are likely to submit statistically significant results in either direction for publication without fear of rejection by journals. Consequently, the resulting published literature is expected to be less prone to publication bias. In such cases, the majority of studies in the meta-analysis yield statistically significant effect estimates in either direction, positioned at the margins of the funnel plot, while small effect estimates from smaller studies are often missing from the insignificant central part of the plot. This phenomenon is referred to as the “tunnel” effect [Peters et al., 2008, Martyn-St James & Carroll, 2010, Onishi & Furukawa, 2014]. The tunnel effect is unlikely to produce a highly biased estimate of the mean effect size, as the funnel plot remains symmetrical, albeit hollow [Doucouliagos et al., 2005, Peters et al., 2008].
It is important to note that funnel plot asymmetry may arise from factors other than publication bias. For example, confounding variables, heterogeneity in true effects, data irregularities, variations in methodological quality, and chance can all contribute to asymmetry [Egger et al., 1997a, Sterne & Egger, 2001, Sterne et al., 2005, 2008, Moreno et al., 2009, Sterne et al., 2011, Callot & Paldam, 2011]. Thus, funnel plot asymmetry cannot be regarded as definitive evidence of publication bias in meta-analysis [Sterne & Egger, 2001]. For this reason, many researchers attribute funnel plot asymmetry to “small-study effects” rather than publication bias, where smaller studies (i.e., those with lower precision) tend to produce larger estimates of a treatment effect compared to larger studies. In other words, smaller studies often report larger effects [Sterne et al., 2000, 2001, Sterne & Egger, 2001, Brockwell & Gordon, 2001]. To address this issue, Peters et al. [2008] propose an enhancement to the traditional funnel plot by incorporating contours that represent conventional “milestone” levels of statistical significance. This “contour-enhanced” funnel plot helps distinguish between asymmetry caused by publication bias and asymmetry attributable to other factors. Briefly, the statistical significance of each point (i.e., effect estimate) on the funnel plot is calculated, and regions of statistical significance are delineated on the plot. If studies appear to be missing from regions of statistical non-significance (or significance), the observed asymmetry is attributed to publication bias (or factors other than publication bias).
The primary limitation of funnel plots is that their visual interpretation, as graphical tools, is inherently subjective. To address this, several statistical methods have been developed to formally test for funnel plot asymmetry. These tests evaluate whether the observed association between the estimated effect size and a measure of sample size (such as the standard error of the effect estimate or its inverse) is greater than what would be expected by chance [Sterne et al., 2008]. Egger et al. [1997a] introduced a widely used regression-based approach to test for funnel plot asymmetry. This method relies on the following specification:
$y_{i}=\beta_{1}+\beta_{0}S_{i}+\varepsilon_{i},$ (23)
where $y_{i}$ represents the estimated empirical effect, and $S_{i}$ denotes its standard error. In the presence of publication selection bias, the reported effect estimates are positively correlated with their standard errors, resulting in an asymmetrical funnel plot. Consequently, a test for funnel plot asymmetry involves evaluating the null hypothesis that the asymmetry coefficient $\beta_{0}$ equals zero. If the hypothesis $H_{0}:\beta_{0}=0$ cannot be rejected, the funnel plot is deemed symmetric [Ashenfelter et al., 1999, Gorg & Strobl, 2001, Sterne & Egger, 2001]. However, the OLS estimator for eq. 23 is often inefficient due to heteroscedasticity, as the standard error of the effect estimates varies widely [Stanley, 2005]. A more suitable approach is weighted least squares (WLS), using the inverse of the variance of the effect estimates ($1/S_{i}^{2}$) as weights. Applying this weighting scheme to eq. 23 is equivalent to performing OLS on the transformed model obtained by dividing both sides of eq. 23 by $S_{i}$:
$\frac{y_{i}}{S_{i}}\equiv t_{i}=\beta_{0}+\beta_{1}\left(\frac{1}{S_{i}}\right)+v_{i},$ (24)
where $t_{i}$ represents the $t$-value for each estimate from the primary studies, $1/S_{i}$ is the precision of the estimate, and $v_{i}=\varepsilon_{i}/S_{i}$ is the transformed error term [Egger et al., 1997a, Sterne et al., 2000, Sterne & Egger, 2005, Stanley, 2005, Stanley et al., 2008]. In this formulation, testing for funnel plot asymmetry is equivalent to testing whether the intercept $\beta_{0}$ equals zero ($H_{0}:\beta_{0}=0$) using a conventional $t$-test. [^7] Although the expected sign of $\beta_{0}$ is typically positive, the possibility of a negative sign necessitates the use of two-sided $p$-values [Sterne & Egger, 2005]. Failure to reject the null hypothesis indicates funnel plot symmetry, suggesting the absence of publication bias. Conversely, rejecting the null implies asymmetry in the funnel plot, indicating the presence of publication bias or, more generally, “small-study” bias. The magnitude and direction of asymmetry are captured by the estimated intercept $\hat{\beta}_{0}$: the greater the deviation of $\hat{\beta}_{0}$ from zero, the more pronounced the asymmetry. If $\hat{\beta}_{0}>0$ ($\hat{\beta}_{0}<0$), the funnel plot is skewed to the right (left). This method is commonly referred to as the “funnel-asymmetry test (FAT)” or the “Egger test” [Egger et al., 1997a, Sterne et al., 2000, Sterne & Egger, 2005, Stanley, 2005, Stanley et al., 2008]. Simulations confirm that the FAT is valid, though it exhibits limited power in detecting publication selection bias [Stanley, 2008]. A practical guideline for interpreting FAT results is as follows: publication selectivity is considered “little to modest” if FAT is not statistically significant or $\lvert\hat{\beta}_{0}\rvert<1$, “substantial” if FAT is significant and $1\leq\lvert\hat{\beta}_{0}\rvert\leq 2$, and “severe” if FAT is significant and $2<\lvert\hat{\beta}_{0}\rvert$ [Stanley & Doucouliagos, 2010, Doucouliagos & Stanley, 2013].
As discussed earlier, there are two sources of publication selection bias: statistical significance and direction of results. The funnel plot is a useful tool for detecting directional selection (type I selection). In cases where there is theoretical support for both significant positive and negative effects, publication selectivity is based on statistical significance. That is, studies with statistically significant effect estimates are more likely to be published, irrespective of their sign (type II selection). In these cases, the funnel plot remains symmetrical but appears hollow and excessively wide. This type of publication selection is generally neglected in meta-analyses because it does not induce a bias in the mean effect size and is thus more benign [Stanley, 2005]. The Galbraith plot [Galbraith, 1988] is a visual tool for identifying the presence of type II publication bias. It is a scatter plot of standardized effect estimates (often a $t$-value) on the vertical axis against the precision of the estimates ($1/S_{i}$) on the horizontal axis. In the absence of type II publication bias, the proportion of $t$-values, $(y_{i}-\mu)/S_{i}$, that fall outside the range of ±$\pm 2$ should not exceed 5% of all cases. Here, ±$\pm 2$ are the approximate critical values at the 5% significance level, and $\mu$ represents the true mean effect, which can be estimated, for instance, using the funnel plot and its top 10% of points that are less prone to selection bias or using PET and PEESE (see below). When there is no true effect beyond publication selection, the $t$-values should be randomly distributed around zero, with no systematic relationship to the precision of the estimates [Stanley, 2005]. In addition to the graphical approach, a statistical test for type II publication bias can be performed by testing the null hypothesis that the slope coefficient in the following regression equals zero ($\beta_{0}=0$):
$\lvert y_{i}\rvert=\beta_{1}+\beta_{0}S_{i}+\varepsilon_{i},$ (25)
where $\lvert y_{i}\rvert$ indicates the absolute value of the $i$th effect estimate and $S_{i}$ is its standard error. As before, this model suffers from heteroscedasticity, rendering the OLS estimator inefficient. This issue can be addressed using WLS with the inverse of the variance of the effect estimates (i.e., $1/S_{i}^{2}$) as the weight, or equivalently, by applying OLS to the transformed model:
$\left\lvert\frac{y_{i}}{S_{i}}\right\rvert\equiv\lvert t_{i}\rvert=\beta_{0}+\beta_{1}\left(\frac{1}{S_{i}}\right)+v_{i}.$ (26)
In the same manner as the FAT, if the null hypothesis $H_{0}:\beta_{0}=0$ is rejected, there is evidence of type II publication selection bias [Stanley, 2005].
Furthermore, a test for the existence of a genuine pooled effect beyond publication bias can be conducted by testing the null hypothesis that the slope coefficient in eq. 24 equals zero ($H_{0}:\beta_{1}=0$). This test is designed to determine whether the true mean effect is statistically different from zero after accounting for publication selection bias. Rejection of the null hypothesis implies the presence of an empirical effect beyond publication bias, with the estimated slope coefficient ($\hat{\beta}_{1}$) indicating the direction and magnitude of the effect. Conversely, failure to reject the null hypothesis suggests no significant evidence of a genuine empirical effect beyond publication bias. This approach is referred to as the “precision-effect test (PET)” [Stanley, 2005, 2008, Sterne & Egger, 2005]. Simulation studies demonstrate that PET is highly effective in detecting publication bias and remains robust regardless of the intensity of the bias [Stanley, 2008]. [^8] Another meta-regression procedure for testing a genuine empirical effect corrected for publication selection relies on a well-established property of statistical power, which states that the absolute value of the standardized test statistic (the $t$-ratio) is positively associated with its degrees of freedom only if a genuine underlying empirical effect exists [Stanley, 2001, 2005]. Accordingly, testing $H_{0}:0\leq\delta_{1}$ in the following regression model provides a formal statistical test, known as the “meta-significance test (MST)”, to detect any genuine empirical effect beyond publication bias: $\text{ln}\lvert t_{i}\rvert=\delta_{0}+\delta_{1}\text{ln}(df_{i})+\varepsilon_{i},$ where ln denotes the natural logarithm, $t_{i}$ is the $t$-value corresponding to the $i$th estimate, and $df_{i}$ represents its degrees of freedom. Rejection of this null hypothesis indicates a statistically significant positive relationship between the absolute value of the test statistic and its degrees of freedom, providing evidence of a genuine nonzero effect [Stanley, 2005].
Although the conventional $t$-test of the null hypothesis $\beta_{1}=0$ in eq. 24 often serves as a valid and powerful test for the existence of a genuine effect beyond publication bias, simulations suggest that $\hat{\beta}_{1}$ tends to underestimate the true mean effect when it is nonzero [Stanley, 2008, Stanley & Doucouliagos, 2014]. To address this issue, the following quadratic approximation is recommended, where the variance of the effect estimate ($S_{i}^{2}$) replaces the standard error ($S_{i}$) in eq. 23, for estimating the genuine pooled effect corrected for publication selection:
$y_{i}=\lambda_{1}+\lambda_{0}S_{i}^{2}+\varepsilon_{i},$ (27)
so that the intercept estimate, $\hat{\lambda}_{1}$, in this formulation serves as a more accurate (less biased) estimate of the true bias-corrected effect. This method is referred to as the “precision-effect estimate with standard error (PEESE)” [Stanley & Doucouliagos, 2007, Moreno et al., 2009, Stanley & Doucouliagos, 2012, 2014]. Given the presence of heteroscedasticity, the WLS estimator can be applied to efficiently estimate the parameters of eq. 27, with weights set to $1/S_{i}^{2}$. Alternatively, this is equivalent to dividing both sides of eq. 27 by $S_{i}$, resulting in:
$\frac{y_{i}}{S_{i}}\equiv t_{i}=\lambda_{0}S_{i}+\lambda_{1}\left(\frac{1}{S_{i}}\right)+v_{i},$ (28)
and subsequently applying OLS to the transformed regression equation. In this case, the coefficient of $1/S_{i}$ in eq. 28 corresponds to the intercept term in eq. 27, whose estimate, $\hat{\lambda}_{1}$, represents the genuine effect corrected for publication bias (PEESE). In summary, if PET results indicate the presence of a true mean effect beyond publication selection (rejecting $H_{0}:\beta_{1}=0$), PEESE, represented by $\hat{\lambda}_{1}$ from eq. 28, should be used. Otherwise, $\hat{\beta}_{1}$ from eq. 24 serves as a less biased estimator of the true bias-corrected effect [Stanley et al., 2017]. This conditional estimator is known as the “precision-effect test and precision-effect estimate with standard error (PET-PEESE)” [Stanley & Doucouliagos, 2014].
To avoid omitted variable bias, the meta-regression model given by eq. 23 can be extended to include a set of moderator variables explaining the heterogeneity in effect sizes ($Z_{k}$’s) and other factors related to the publication selection process ($K_{j}$’s), as follows:
$y_{i}=\beta_{1}+\sum_{k}\alpha_{k}Z_{ki}+\beta_{0}S_{i}+\sum_{j}\gamma_{j}K_{ji}S_{i}+\varepsilon_{i},$ (29)
where $\beta_{1}$ and $\beta_{0}S_{i}$ in eq. 23 are replaced by $\beta_{1}+\sum_{k}\alpha_{k}Z_{ki}$ and $\beta_{0}S_{i}+\sum_{j}\gamma_{j}K_{ji}S_{i}$, respectively [Stanley et al., 2008, Doucouliagos & Stanley, 2009, Stanley & Doucouliagos, 2012]. Due to the presence of heteroscedasticity, this model can be estimated using WLS with $1/S_{i}^{2}$ as weights or, equivalently, by applying OLS to the transformed model.  [^9] Similarly, the extended versions of eqs. 27 and 28 can be expressed as: $y_{i}=\lambda_{1}+\sum_{k}\alpha_{k}Z_{ki}+\lambda_{0}S_{i}^{2}+\sum_{j}\gamma_{j}K_{ji}S_{i}^{2}+\varepsilon_{i},$ and $\frac{y_{i}}{S_{i}}\equiv t_{i}=\lambda_{0}S_{i}+\sum_{j}\gamma_{j}K_{ji}S_{i}+\lambda_{1}\left(\frac{1}{S_{i}}\right)+\sum_{k}\alpha_{k}\left(\frac{Z_{ki}}{S_{i}}\right)+v_{i},$ respectively [Doucouliagos et al., 2012, Stanley & Doucouliagos, 2014]. These models can be used to identify factors explaining the heterogeneity in effect sizes. While eqs. 27 and 28 are preferable to eqs. 23 and 24 for estimating the true mean effect beyond publication selection, it remains unclear whether the multiple meta-regression models described above are superior to eqs. 29 and 30 for explaining variation in true effects [Doucouliagos et al., 2012]. Similarly, the transformed version of eq. 29 can be written as:
$\frac{y_{i}}{S_{i}}\equiv t_{i}=\beta_{0}+\sum_{j}\gamma_{j}K_{ji}+\beta_{1}\left(\frac{1}{S_{i}}\right)+\sum_{k}\alpha_{k}\left(\frac{Z_{ki}}{S_{i}}\right)+v_{i}.$ (30)
The coefficients $\gamma_{j}$ represent the effects of the corresponding variables ($K_{j}$’s) on publication selection, while the coefficients $\alpha_{k}$ capture the effects of the corresponding moderator variables ($Z_{k}$’s) on the true effect size, after accounting for publication selection. A Wald test can be conducted to evaluate the joint hypothesis that $\beta_{0}$ and $\gamma_{j}$’s are equal to zero. Rejection of this null hypothesis indicates the presence of publication selection bias. Similarly, rejecting the joint hypothesis that the $\alpha_{k}$ coefficients are zero suggests that the $Z$-variables account for at least part of the heterogeneity in the true effects.
[…]
5 Unrestricted Weighted Least Squares (UWLS) Estimator for Meta-Analysis and Meta-Regression Models
As discussed so far, in the context of meta-analysis and meta-regression analysis, random effects models account for between-study heterogeneity in true effect sizes. A random effects meta-analysis model incorporates an additive between-study variance component to capture the heterogeneity in true effect sizes, beyond the sampling variability. Also, in random effects meta-regression model, moderators are introduced to explain some of the variability in true effect sizes, but the model still incorporates an additive variance component to account for residual heterogeneity that remains unexplained by the moderators. In contrast, a fixed effect meta-analysis model does not allow for any between-study heterogeneity, and a fixed effects meta-regression model does not account for residual heterogeneity. Some authors (e.g., Thompson & Sharp, 1999, Higgins & Thompson, 2004) caution against using a fixed effects meta-regression. They argue that, in practice, the included moderators often do not account for all the between-study heterogeneity. This can lead to residual heterogeneity that is not captured by the model, resulting in coefficient standard errors that are generally too small. Therefore, relying on conventional fixed effects meta-analysis and meta-regression models can lead to misleading conclusions, as they may ignore all or part of the variability in true effect sizes across studies. To address this issue, one approach is to incorporate total heterogeneity into the fixed effect meta-analysis model or residual heterogeneity into fixed effects meta-regression model by applying a common multiplicative factor, $\sigma^{2}$ (which is denoted as the overdispersion parameter $\phi$ in some studies), to each of the variances $\sigma_{i}^{2}$, resulting in $\sigma^{2}\sigma_{i}^{2}$ or $\phi\sigma_{i}^{2}$. This is equivalent to estimating fixed effects models using unrestricted weighted least squares (UWLS) instead of the restricted weighted least squares (WLS) estimator ($\sigma^{2}=1$) that was used in the previous sections [Thompson & Sharp, 1999, Stanley & Doucouliagos, 2015, 2016]. In the following discussion, the UWLS estimator for the fixed effect meta-analysis and fixed effects meta-regression models will be explained.
The fixed effect meta-analysis model can be rewritten as follows:
$\bm{y}=\mu_{F}\bm{1}+\bm{\varepsilon},\quad\bm{\varepsilon}\sim N(\bm{0},\bm{\Sigma}),$ (31)
where $\bm{1}$ is a $k$-dimensional vector of ones, $\bm{y}=(y_{1},y_{2},\dots,y_{k})^{\top}$ represents the column vector of observed effect sizes, and $\bm{\varepsilon}=(\varepsilon_{1},\varepsilon_{2},\dots,\varepsilon_{k})^{\top}$ is the vector of errors. Additionally, $E(\bm{\varepsilon})=\bm{0}$, and $Var(\bm{\varepsilon})=E(\bm{\varepsilon}\bm{\varepsilon}^{\prime})=\sigma^{2}\bm{\Omega}=\bm{\Sigma}$. For heteroscedastic but uncorrelated error terms, the error variance-covariance matrix ($\bm{\Omega}$) is diagonal:
$\bm{\Omega}=\begin{pmatrix}\sigma_{1}^{2}&0&\cdots&0\\ 0&\sigma_{2}^{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\sigma_{k}^{2}\end{pmatrix},$
with the inverse matrix given by:
$\bm{\Omega}^{-1}=\begin{pmatrix}\frac{1}{\sigma_{1}^{2}}&0&\cdots&0\\ 0&\frac{1}{\sigma_{2}^{2}}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\frac{1}{\sigma_{k}^{2}}\end{pmatrix}=\underbrace{\begin{pmatrix}w_{1}&0&\cdots&0\\ 0&w_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&w_{k}\end{pmatrix}}_{\bm{W}}$ (32)
and $w_{i}=\frac{1}{\sigma_{i}^{2}}$ for $i=1,2,\dots,k$. Then, the inverse of $\bm{\Sigma}$ is:
$\bm{\Sigma}^{-1}=\begin{pmatrix}\frac{1}{\sigma^{2}\sigma_{1}^{2}}&0&\cdots&0\\ 0&\frac{1}{\sigma^{2}\sigma_{2}^{2}}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\frac{1}{\sigma^{2}\sigma_{k}^{2}}\end{pmatrix}=\underbrace{\begin{pmatrix}w_{1}^{u}&0&\cdots&0\\ 0&w_{2}^{u}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&w_{k}^{u}\end{pmatrix}}_{\bm{W^{u}}}$ (33)
with $w_{i}^{u}=\frac{1}{\sigma^{2}\sigma_{i}^{2}}=\frac{1}{\sigma^{2}}w_{i}$ for $i=1,2,\dots,k$.
In practice, $\sigma^{2}_{i}$’s are estimated by the sample variances obtained from primary studies ($S^{2}_{i}$’s), although they are conventionally treated as fixed and known in the analysis. In addition, $\sigma^{2}>0$ as a multiplicative variance parameter can be routinely estimated using the mean squared error (MSE) of the estimated intercept-only regression model by the WLS estimator as follows:
$s^{2}=\frac{\hat{\bm{\varepsilon}}^{*^{\prime}}\hat{\bm{\varepsilon}}^{*}}{k-1}=\frac{(\bm{y}-\hat{\mu}_{F}\bm{1})^{\prime}\hat{\bm{W}}(\bm{y}-\hat{\mu}_{F}\bm{1})}{k-1}$ (34)
where $\hat{\bm{\varepsilon}}^{*}=\bm{y}^{*}-\hat{\mu}_{F}\bm{1}$, $\bm{y}^{*}=\hat{\bm{W}}^{1/2}\bm{y}$, and $\hat{\bm{W}}=\text{diag}[\hat{w}_{1},\dots,\hat{w}_{k}]=\text{diag}\left[\frac{1}{S_{1}^{2}},\dots,\frac{1}{S_{k}^{2}}\right]$. Therefore, $\hat{w}_{u_{i}}=\frac{1}{S^{2}S_{i}^{2}}$ serves as an estimate of the “true” weight $w_{i}^{u}=\frac{1}{\sigma^{2}\sigma_{i}^{2}}$.
Then, the UWLS estimator for the meta-analysis model is given by:
$\hat{\mu}_{U}=\frac{\sum_{i=1}^{k}\hat{w}_{i}^{u}y_{i}}{\sum_{i=1}^{k}\hat{w}_{i}^{u}}=\frac{\sum_{i=1}^{k}\frac{1}{S^{2}S_{i}^{2}}y_{i}}{\sum_{i=1}^{k}\frac{1}{S^{2}S_{i}^{2}}}=\frac{\sum_{i=1}^{k}\frac{1}{S_{i}^{2}}y_{i}}{\sum_{i=1}^{k}\frac{1}{S_{i}^{2}}}=\frac{\sum_{i=1}^{k}\hat{w}_{i}y_{i}}{\sum_{i=1}^{k}\hat{w}_{i}}=\hat{\mu}_{F},$ (35)
whose variance is estimated as:
$\widehat{Var}(\hat{\mu}_{U})=\frac{1}{\sum_{i=1}^{k}\hat{w}_{i}^{u}}=\frac{S^{2}}{\sum_{i=1}^{k}\hat{w}_{i}}=S^{2}\widehat{Var}(\hat{\mu}_{F}).$ (36)
Consequently, the WLS and UWLS estimators provide identical point estimates for the fixed effect meta-analysis model (i.e., $\hat{\mu}_{U}=\hat{\mu}_{F}$), but their estimated variances differ only by a multiplicative factor of $S^{2}$. In fact, WLS is a special case of UWLS, with the constraint $\sigma^{2}=1$ imposed on it.
Similarly, the fixed effects meta-regression model can be expressed as:
$\bm{y}=\bm{X}\bm{\beta}_{F}+\bm{\varepsilon},\quad\bm{\varepsilon}\sim N(\bm{0},\bm{\Sigma}),$ (37)
where $\bm{X}$ is the $k\times(p+1)$ design matrix containing the moderators (independent variables) as columns, with the first column consisting of ones to account for the intercept term. $\bm{\beta}_{F}$ is a $(p+1)\times 1$ vector of fixed effect coefficients to be estimated.
The UWLS estimator for this model is:
$\hat{\bm{\beta}}_{U}=\left[\bm{X}^{\prime}\bm{\Sigma}^{-1}\bm{X}\right]^{-1}\left[\bm{X}^{\prime}\bm{\Sigma}^{-1}\bm{y}\right]=\left[\bm{X}^{\prime}(\sigma^{2}\bm{\Omega})^{-1}\bm{X}\right]^{-1}\left[\bm{X}^{\prime}(\sigma^{2}\bm{\Omega})^{-1}\bm{y}\right].$
After canceling out the term $\sigma^{2}$, this becomes:
$\hat{\bm{\beta}}_{U}=\left(\bm{X}^{\prime}\bm{\Omega}^{-1}\bm{X}\right)^{-1}\left(\bm{X}^{\prime}\bm{\Omega}^{-1}\bm{y}\right)=\left(\bm{X}^{\prime}\bm{W}\bm{X}\right)^{-1}\left(\bm{X}^{\prime}\bm{W}\bm{y}\right).$
Then, replacing $\bm{W}=\text{diag}[w_{1},\dots,w_{k}]$ with $\hat{\bm{W}}=\text{diag}[\hat{w}_{1},\dots,\hat{w}_{k}]$ results in the following:
$\hat{\bm{\beta}}_{U}=\left(\bm{X}^{\prime}\hat{\bm{W}}\bm{X}\right)^{-1}\left(\bm{X}^{\prime}\hat{\bm{W}}\bm{y}\right)=\hat{\bm{\beta}}_{F},$ (38)
with the estimated variance-covariance matrix:
$\widehat{Var}(\hat{\bm{\beta}}_{U})=\left(\bm{X}^{\prime}\hat{\bm{W}}^{U}\bm{X}\right)^{-1}=S^{2}\left(\bm{X}^{\prime}\hat{\bm{W}}\bm{X}\right)^{-1}=S^{2}\widehat{Var}(\hat{\bm{\beta}}_{F}),$ (39)
where $S^{2}$ is an estimator of $\sigma^{2}$, computed as the mean squared error (MSE) of the estimated meta-regression model by the WLS estimator, as follows:
$S^{2}=\frac{\hat{\bm{\varepsilon}}^{*^{\prime}}\hat{\bm{\varepsilon}}^{*}}{k-p-1}=\frac{\left(\bm{y}-\bm{X}\hat{\bm{\beta}}_{F}\right)^{\prime}\hat{\bm{W}}\left(\bm{y}-\bm{X}\hat{\bm{\beta}}_{F}\right)}{k-p-1},$ (40)
where $\hat{\bm{\varepsilon}}^{*}=\bm{y}^{*}-\bm{X}^{*}\hat{\bm{\beta}}_{F}$, $\bm{y}^{*}=\hat{\bm{W}}^{1/2}\bm{y}$, and $\bm{X}^{*}=\hat{\bm{W}}^{1/2}\bm{X}$.
Consequently, the WLS and UWLS estimators provide identical point estimates for the fixed effects meta-regression model (i.e., $\hat{\bm{\beta}}_{U}=\hat{\bm{\beta}}_{F}$), but their estimated variances differ only by a multiplicative factor of $S^{2}$.
Overall, both conventional random effects models and multiplicative fixed effects models account for between-study heterogeneity. However, Stanley & Doucouliagos [2015, 2016] demonstrate that in both meta-analysis and meta-regression analysis, a multiplicative fixed effects (UWLS) estimator outperforms a conventional random effects estimator (i.e., has smaller bias and MSE) when there is publication selection bias or selective reporting bias. Even when the random effects model is known to be true and there is no publication selection bias, simulations show that the random effects model has no significant advantage over UWLS [Stanley et al., 2017]. Further supporting this, Stanley et al. [2023] conducted a comprehensive analysis of 67,308 meta-analyses from the Cochrane Database of Systematic Reviews (CDSR) published between 1997 and 2020. They compared UWLS with the conventional random effects model to evaluate which better represents medical research. Their findings reveal that UWLS frequently outperforms the random effects model, regardless of the level of heterogeneity, the number of studies, and the outcome measures. The authors conclude that UWLS is generally a superior model for medical research and should be routinely reported in all meta-analyses.
6 Effect Size Dependence
So far, it has been assumed that all effect size estimates are statistically independent of one another. This assumption holds when all studies are independent and each study contributes only one effect estimate to the meta-analysis [Hedges et al., 2010]. In practice, however, it is quite common for effect size estimates to be dependent for various reasons. According to Hedges et al. [2010], dependence among effect size estimates across different studies may occur either through the estimation errors ($\varepsilon_{i}$’s in eq. 1), the effect size parameters ($\beta_{i}$’s in eq. 1), or both. Dependence among the estimation errors arises when a primary study reports more than one effect size estimate—i.e., multiple effect sizes are clustered (or nested) within one or more studies, which is referred to as “correlated effects.” This situation arises, for example, when effect size estimates share a common control group, when multiple outcomes are measured for the same individuals, or when the same outcome is measured at multiple time points for the same individuals [Becker, 2000, Hedges et al., 2010, Cheung, 2015a, Tipton, 2015, López-López et al., 2017]. This type of dependence is always present in the data, even when a fixed effect model, where the true effect sizes are fixed, is employed [Cheung, 2015a]. Dependence among the (random) effect parameters occurs when there are groups or clusters of studies that share common characteristics, such as those conducted by the same investigator/research group or in the same laboratory, country, or region (i.e., multiple studies are nested within larger clusters). This type of dependence is referred to as “hierarchical effects” or “hierarchical dependence” [Stevens & Taylor, 2009, Hedges et al., 2010].
Ignoring or misspecifying dependence in meta-analytic data can lead to invalid estimates of the mean effect size (or the meta-regression coefficients) and their standard errors, resulting in inflated Type I error rates [Becker, 2000, Olkin & Gleser, 2009]. Several strategies have been employed to address statistically dependent effect sizes. The simplest of these involve ignoring the dependence among effect size estimates, selecting only one effect size estimate per study, computing a single summary effect size per study by averaging the effect size estimates clustered within studies, and shifting the unit of analysis. Although these strategies can effectively remove the dependence among effect size estimates, they result in a loss of information (see Rosenthal & Rubin, 1986, Cooper, 1998, Marín-Martínez & Sánchez-Meca, 1999, Becker, 2000, Jackson et al., 2011, Scammacca et al., 2014, Cheung, 2014, 2015a, Borenstein et al., 2021). A more principled approach is to explicitly model the correlations among effect size estimates using multivariate meta-analytic methods. The challenge with this approach is that it generally requires knowledge of the covariance structure of within-study effect size estimates, which is rarely reported in primary studies (see Raudenbush et al., 1988, Kalaian & Raudenbush, 1996, Becker, 2000, Olkin & Gleser, 2009, Jackson et al., 2011, Mavridis & Salanti, 2013). Fortunately, two alternative approaches—the robust variance estimation (RVE) and the three-level meta-analysis model—circumvent this issue. These recently developed approaches are discussed in the following subsections.
6.1 Robust Variance Estimation (RVE)
Hedges et al. [2010] propose an appealing meta-analytic method, known as robust variance estimation (RVE), which can be used to handle both types of clustering or dependence structures in meta-analysis (i.e., correlated effects and hierarchical effects) without requiring knowledge of the covariance structure of the effect estimates and without making strong assumptions about the error distribution. Instead of modeling dependence as done in multivariate meta-analytic methods, the RVE method adjusts the variance-covariance matrix of meta-regression coefficients to account for the dependence of effect sizes, thereby improving statistical inference, including confidence intervals and hypothesis testing. This method is briefly described in the following discussion. For further details, the interested reader is referred to Hedges et al. [2010], Tipton [2013], Tanner-Smith & Tipton [2014], Tipton [2015], Tipton & Pustejovsky [2015], Tanner-Smith et al. [2016].
Consider the general case of a meta-analysis of $m$ studies, with $k_{j}$ effect size estimates nested within the $j$-th study ($j=1,\dots,m$), such that the total number of effect size estimates is $\sum_{j=1}^{m}k_{j}=k$. Then, a general meta-regression model can be written as
$\bm{y}=\bm{X}\bm{\beta}+\bm{\epsilon},$ (41)
where $\bm{y}=(\bm{y}_{1}^{{}^{\prime}},\dots,\bm{y}_{m}^{{}^{\prime}})^{\prime}$ is a $k\times 1$ vector of $m$ vectors of effect size estimates, each with $k_{j}$ elements; $\bm{\epsilon}=(\bm{\epsilon}_{1}^{{}^{\prime}},\dots,\bm{\epsilon}_{m}^{{}^{\prime}})^{\prime}$ is a stack of $m$ error vectors, each with mean $\bm{0}$ and variance-covariance matrix $\bm{\Sigma}_{j}$; $\bm{X}=(\bm{X}_{1}^{{}^{\prime}},\dots,\bm{X}_{m}^{{}^{\prime}})^{\prime}$ is a stack of $k_{j}\times(p+1)$ matrices of 1s in the first column and observations on $p$ covariates in the other columns (note that the covariates may vary across effect sizes or studies); and $\bm{\beta}=(\beta_{0},\beta_{1},\dots,\beta_{p})^{\prime}$ is a $(p+1)\times 1$ vector of parameters, which can be estimated by weighted least squares as
$\bm{\hat{\beta}}=\left(\sum_{j=1}^{m}\bm{X}_{j}^{\prime}\bm{W}_{j}\bm{X}_{j}\right)^{-1}\left(\sum_{j=1}^{m}\bm{X}_{j}^{\prime}\bm{W}_{j}\bm{y}_{j}\right),$ (42)
where $\bm{W}_{j}$ is the weight matrix associated with the $j$-th study. The exact variance-covariance matrix of $\bm{\hat{\beta}}$ can be written as
$Var(\bm{\hat{\beta}})=\left(\sum_{j=1}^{m}\bm{X}_{j}^{\prime}\bm{W}_{j}\bm{X}_{j}\right)^{-1}\left(\sum_{j=1}^{m}\bm{X}_{j}^{\prime}\bm{W}_{j}\bm{\Sigma}_{j}\bm{W}_{j}\bm{X}_{j}\right)\left(\sum_{j=1}^{m}\bm{X}_{j}^{\prime}\bm{W}_{j}\bm{X}_{j}\right)^{-1}.$ (43)
The problem with this estimator is that the within-study variance-covariance matrices, $\bm{\Sigma}_{1},\dots,\bm{\Sigma}_{m}$, are usually unknown and need to be estimated. In the standard fixed and random effects models, as discussed earlier, the variance elements are assumed to be fully known (in the fixed effect case) or partially known and partially estimated (in the random effects case). However, the covariance elements are assumed to be equal to zero. In practice, this assumption usually does not hold, especially when some studies contribute more than one effect size estimate to the meta-analysis. The complications arise when the covariance structure of the effect size estimates is unknown, which is the case in most applications. In the RVE method, the cross-products of observed residuals within study $j$ are used to estimate the elements of $\bm{\Sigma}_{j}$. Thus, the RVE estimator can be written as
$\widehat{Var}(\bm{\hat{\beta}})=\left(\sum_{j=1}^{m}\bm{X}_{j}^{\prime}\bm{W}_{j}\bm{X}_{j}\right)^{-1}\left(\sum_{j=1}^{m}\bm{X}_{j}^{\prime}\bm{W}_{j}\bm{\hat{\epsilon}}_{j}\bm{\hat{\epsilon}}_{j}^{\prime}\bm{W}_{j}\bm{X}_{j}\right)\left(\sum_{j=1}^{m}\bm{X}_{j}^{\prime}\bm{W}_{j}\bm{X}_{j}\right)^{-1},$ (44)
where $\bm{\hat{\epsilon}}_{j}=\bm{y}_{j}-\bm{X}_{j}\bm{\hat{\beta}}$ is a $k_{j}\times 1$ vector containing all $k_{j}$ residuals for the $k_{j}$ effect size estimates nested within study $j$. According to Hedges et al. [2010], although the $\bm{\hat{\Sigma}}_{j}=\bm{\hat{\epsilon}}_{j}\bm{\hat{\epsilon}}_{j}^{\prime}$ are rather poor estimates of the $\bm{\Sigma}_{j}$’s, $\widehat{Var}(\bm{\hat{\beta}})$ will converge in probability to the true covariance matrix, $Var(\bm{\hat{\beta}})$, if $m\to\infty$.
An important property of the RVE estimator is that it provides (asymptotically) accurate estimates of the standard errors and a useful framework for statistical inference, regardless of the weights used. That is, RVE is valid for any given choice of $\bm{W}_{j}$’s. As noted by Hedges et al. [2010], this means that the only criterion for choosing the weighting scheme is efficiency. In this case, the “optimal” weighting matrices, that is, the $\bm{W}_{j}$’s that produce the most efficient estimator of $\bm{\beta}$, would be $\bm{W}_{j}=\bm{\Sigma}_{j}^{-1}$. Unfortunately, as discussed above, the true covariance structure of the effect size estimates is usually unknown (thus necessitating a robust method). For this reason, Hedges et al. [2010] suggest approximating the optimal weights based on a “working” model of the unknown covariance structure, which produces an approximately efficient estimator of $\bm{\beta}$. The resulting approximately optimal weighting matrix for the $j$-th study is defined as $\bm{W}_{j}\approx\bm{\Sigma}_{aj}^{-1}$, where $\bm{\Sigma}_{aj}$ is the working covariance matrix. They propose two working models for handling and coping with two types of dependence structures (i.e., correlated and hierarchical effects) in meta-analytic research separately (see Hedges et al., 2010, Tipton, 2015, Tipton & Pustejovsky, 2015).
In the correlated effects case, the working model can be written as:
$\bm{\Sigma}_{aj}=\tau^{2}\bm{J}_{j}+\rho\sigma_{j}^{2}(\bm{J}_{j}-\bm{I}_{j})+\sigma_{j}^{2}\bm{I}_{j},$ (45)
where $\bm{J}_{j}$ is a $k_{j}\times k_{j}$ matrix of ones, $\bm{I}_{j}$ is a $k_{j}\times k_{j}$ identity matrix, and $\tau^{2}$ represents the between-study variance as a measure of heterogeneity in the study-average true effect sizes. In this model, the sampling variances associated with multiple effect sizes nested within each study are assumed to be equal ($\sigma_{ij}^{2}=\sigma_{j}^{2}$), which is the case when the sample sizes are equal. Moreover, the simplified working model assumes a common correlation ($\rho$) for all pairs of effect size estimates. Based on this working model of the known covariance structure, Hedges et al. [2010] propose the approximately inverse variance (optimal) weights for study $j$ as follows:
$\bm{W}_{j}=w_{j}\bm{I}_{j}=\left\{1/[k_{j}(\sigma_{j}^{2}+\tau^{2})]\right\}\bm{I}_{j}.$ (46)
In other words, $\bm{W}_{j}=diag[w_{1j},...,w_{k_{j}j}]$, where $w_{ij}=w_{j}=1/k_{j}(\sigma_{j}^{2}+\tau^{2})$. Hedges et al. [2010] provide a method of moments (MM) estimator of $\tau^{2}$, depending on the unknown value of $\rho\in(0,1)$. They show that in many practical situations, the choice of $\rho$ plays a very minor role in the estimation of $\tau^{2}$. However, they recommend performing a sensitivity analysis to ensure that $\hat{\tau}^{2}$ is not highly sensitive to the choice of $\rho$. In addition, the parameter $\sigma_{j}^{2}$ can be estimated by the average of within-study sampling (error) variances, $S_{j}^{2}=\sum_{i=1}^{k_{j}}\frac{1}{k_{j}}S_{ij}^{2}$. Note that the fixed effects weights can be obtained by setting $\tau^{2}$ equal to zero.
In the hierarchical effects case, the working model can be written as
$\bm{\Sigma}_{aj}=\tau^{2}\bm{J}_{j}+\omega^{2}\bm{I}_{j}+\bm{\sigma}_{j}^{2},$ (47)
where $\bm{J}_{j}$ is a $k_{j}\times k_{j}$ matrix of ones, $\bm{I}_{j}$ is a $k_{j}\times k_{j}$ identity matrix, $\tau^{2}$ is the between-study variance as a measure of heterogeneity in the study-average true effects (as before), $\omega^{2}$ is the within-study variance as a measure of heterogeneity in the within-study true effects, and $\bm{\sigma}_{j}^{2}=diag[\sigma_{1j}^{2},...,\sigma_{k_{j}j}^{2}]$ is a $k_{j}\times k_{j}$ diagonal matrix whose diagonal elements are the sampling (error) variances in study $j$. Based on this working model, Hedges et al. [2010] propose the approximately inverse variance weights for study $j$ as follows:
$\bm{W}_{j}=diag[w_{1j},...,w_{k_{j}j}],\qquad\text{where}\qquad w_{ij}=1/(\sigma_{ij}^{2}+\omega^{2}+\tau^{2}).$ (48)
Hedges et al. [2010] develop the method of moments (MM) estimators for the parameters $\omega^{2}$ and $\tau^{2}$. The $\sigma_{ij}^{2}$ are estimated by the sample variances $S_{ij}^{2}$ in practice, while they are commonly considered to be fixed and known in the analysis.
In practice, it is common for both types of dependence to occur in the same meta-analysis. In this case, the choice between the two working models can be based on the most common type of dependence in the data to be meta-analyzed, as recommended by Tanner-Smith & Tipton [2014] and Tipton & Pustejovsky [2015].
The RVE estimator can be used to construct Wald-type test statistics for testing hypotheses or forming confidence intervals about one or more elements of $\bm{\beta}$ (see Hedges et al., 2010, Tipton & Pustejovsky, 2015). For example, a $100(1-\alpha)\%$ confidence interval for the parameter $\beta_{k}$ ($k=0,1,\dots,p$) can be calculated as $\hat{\beta}_{k}\pm t_{m-p-1,\alpha/2}\sqrt{\widehat{Var}(\hat{\beta}_{k})}$, where $\widehat{Var}(\hat{\beta}_{k})$ is the $k+1$th diagonal element of $\widehat{Var}(\bm{\hat{\beta}})$, and $t_{m-p-1,\alpha/2}$ is the upper $\alpha/2$ quantile of the $t$-distribution with $m-p-1$ degrees of freedom. However, when the sample size is small or moderate, simulation studies suggest that tests based on the RVE estimator may produce inflated Type I error rates [Hedges et al., 2010, Tipton, 2013, 2015, Tipton & Pustejovsky, 2015]. Therefore, several small-sample adjustments have been proposed by Hedges et al. [2010], Tipton [2015], Tipton & Pustejovsky [2015]. These adjustments fall into two categories: one for the residuals used in the RVE estimator, given by
$Adj.\widehat{Var}(\bm{\hat{\beta}})=\left(\sum_{j=1}^{m}\bm{X}_{j}^{\prime}\bm{W}_{j}\bm{X}_{j}\right)^{-1}\left(\sum_{j=1}^{m}\bm{X}_{j}^{\prime}\bm{W}_{j}\bm{A}_{j}\bm{\hat{\epsilon}}_{j}\bm{\hat{\epsilon}}_{j}^{\prime}\bm{A}_{j}\bm{W}_{j}\bm{X}_{j}\right)\left(\sum_{j=1}^{m}\bm{X}_{j}^{\prime}\bm{W}_{j}\bm{X}_{j}\right)^{-1},$ (49)
where $\bm{A}_{j}$ is a $k_{j}\times k_{j}$ adjustment matrix. The other adjustment concerns the degrees of freedom associated with the distribution of Wald-type test statistics ($t$, $F$, and $\chi^{2}$) used for constructing confidence intervals or testing hypotheses about single and multiparameter meta-regression coefficients (and the overall mean effect).
RVE can be implemented in R (via the ‘$robumeta$’ package, Fisher & Tipton, 2015, Tanner-Smith et al., 2016, Fisher et al., 2017), Stata (via the ‘$ROBUMETA$’ macro, Hedberg, 2011, Tanner-Smith & Tipton, 2014), and, to some extent, SPSS (see Tanner-Smith & Tipton, 2014).
6.2 Three-Level Model
An alternative way to handle the hierarchical dependence of the effect sizes is to use a multilevel (three-level) meta-analysis model (see Hox et al., 2010, Konstantopoulos, 2011, Van den Noortgate et al., 2013, 2015, Cheung, 2014, 2015a). It is evident that the random effects meta-analysis model in eq. 4 has a two-level structure as follows:
$\begin{split}\text{Level\;1}:\qquad y_{i}&=\beta_{i}+\varepsilon_{i},\qquad\qquad\;\;\;\;\;\varepsilon_{i}\sim N(0,\sigma_{i}^{2}),\\ \text{Level\;2}:\qquad\beta_{i}&=\mu_{R}+u_{i},\qquad\qquad\;\;\;u_{i}\sim N(0,\tau^{2}),\\ \text{Combined}:\;\;\;y_{i}&=\mu_{R}+u_{i}+\varepsilon_{i}.\qquad\qquad\qquad\qquad\quad\end{split}$ (50)
Similarly, the random/mixed effects meta-regression model in eq. 17 can be expressed as a two-level model (simply by adding the moderator variables to the second level of the previous model). The two-level random effects meta-analysis model can be extended to a three-level model by adding another level to account for the clustering of effect sizes within studies, as follows:
$\begin{split}\text{Level\;1}:\qquad y_{ij}=\beta_{ij}+\varepsilon_{ij},\qquad\qquad\varepsilon_{ij}\sim N(0,\sigma_{ij}^{2}),\\ \text{Level\;2}:\qquad\beta_{ij}=\theta_{j}+\upsilon_{ij},\qquad\qquad\;\upsilon_{ij}\sim N(0,\omega^{2}),\\ \text{Level\;3}:\qquad\theta_{j}=\mu_{R}+u_{j},\qquad\qquad\;\;u_{ij}\sim N(0,\tau^{2}),\end{split}$ (51)
where $y_{ij}$ is an estimate of the $i$th true (population) effect size in the $j$th study ($\beta_{ij}$), $\theta_{j}$ is the average true effect in the $j$th study, and $\mu_{R}$ is the overall average true effect (i.e., the average of all true effect sizes). Substituting the level 3 equation into the level 2 equation, and then substituting the resulting equation into the level 1 equation yields the following combined equation:
$y_{ij}=\mu_{R}+u_{j}+\upsilon_{ij}+\varepsilon_{ij},\qquad\qquad\qquad\;$ (52)
where the random effects at different hierarchical levels ($\upsilon_{ij}$ and $u_{j}$) and the sampling error ($\varepsilon_{ij}$) are assumed to be independent of each other, i.e., $Cov(\upsilon_{ij},u_{j})=Cov(\upsilon_{ij},\varepsilon_{ij})=Cov(u_{j},\varepsilon_{ij})=0$. In this formulation, there are three potential sources of variation in effect size estimates: $\varepsilon_{ij}$ and $u_{j}$ as in the traditional two-level meta-analytic model, and $\upsilon_{ij}$ as an additional source of random variation. Thus, the variance of the $ij$th effect size estimate is given by $Var(y_{ij})=\sigma_{ij}^{2}+\omega^{2}+\tau^{2}$, where $Var(y_{ij})$ is the variance of the $ij$th effect size estimate, $\sigma_{ij}^{2}$ is the sampling (error) variance associated with the $ij$th effect size estimate (which is considered to be known), $\omega^{2}$ is the within-study variance, representing the heterogeneity in the within-study true effect sizes, and $\tau^{2}$ is the between-study variance, representing the heterogeneity in the study-average true effect sizes. That is, the total variation in effect size estimates is the sum of the level 2 and level 3 heterogeneity, along with the sampling (error) variance. Although the multilevel (three-level) model proposed for dealing with multiple outcomes within the same study is based on the assumption that there is no sampling covariation (i.e., independent samples at the first level), a simulation study by Van den Noortgate et al. [2013] suggests that adding an intermediate level of outcomes within studies plays an important role in accurately accounting for the sampling covariance, providing appropriate standard errors and offering a useful framework for statistical inference about the effects.
In a similar fashion to the traditional two-level random effects meta-analysis, the first step in this extended meta-analysis is to estimate the model parameters (i.e., the true mean effect, $\mu_{R}$, and the variance components of the random effects, $\omega^{2}$ and $\tau^{2}$). The usual estimators of the parameters of multilevel regression models are maximum likelihood (ML) and restricted maximum likelihood (REML), which estimate all parameters simultaneously [Hox et al., 2010, Konstantopoulos, 2011, Van den Noortgate et al., 2013, 2015, Cheung, 2014, 2015a]. In general, the ML and REML procedures produce very similar estimates of variance components. However, if they do not, the REML procedure is preferred because it accounts for the loss of degrees of freedom due to the estimation of regression coefficients as fixed effects, and thus provides less biased estimates of variance components [Hox et al., 2010]. As in the case of traditional two-level random effects meta-analysis, the Hartung-Knapp-Sidik-Jonkman method can be used to test hypotheses or form confidence intervals about the mean effect size, $\mu_{R}$, instead of the Wald test based on the $z$ distribution.
The next step involves testing for the heterogeneity of the true effects based on Cochran’s $Q$ statistic. When the true effect sizes are homogeneous, cluster effects do not exist, and thus the conventional $Q$ statistic given by eq. 7 can be applied directly. Under the null hypothesis of homogeneity, this statistic is asymptotically distributed as $\chi^{2}$ with $k-1$ degrees of freedom, where $k=\sum_{j=1}^{m}k_{j}$ is the total number of effect sizes. Large values of $Q$ provide evidence against the validity of the null hypothesis. The rejection of the null hypothesis, however, is not very informative, as the total heterogeneity in the true effects is the sum of two components: within-study heterogeneity at level 2 and between-study heterogeneity at level 3. Therefore, one might be interested in testing the statistical significance of the two heterogeneity parameters separately. This can be achieved by using the likelihood ratio (LR) or chi-square difference test, which is based on the likelihood function [Van den Noortgate et al., 2015, Cheung, 2014, 2015a]. The null hypotheses to be tested are: $\text{H}_{0}:\omega^{2}=0$ and $\text{H}_{0}:\tau^{2}=0$. If the null hypothesis of $\omega^{2}=0$ is rejected, there is within-study heterogeneity in the true effects. Similarly, the rejection of the null hypothesis of $\tau^{2}=0$ indicates the presence of between-study heterogeneity. However, there is a problem with this testing procedure. The hypothesis that $\omega^{2}=0$ or $\tau^{2}=0$ is not well-defined for the LR test because the true parameter under the null hypothesis lies on the boundary of the parameter space ($\omega^{2}\geq 0$ and $\tau^{2}\geq 0$). In this situation, the test statistic does not follow a $\chi^{2}$ distribution with one degree of freedom, even when the null hypothesis is true [Stoel et al., 2006, Hox et al., 2010, Greene, 2012, Cheung, 2014, 2015a]. One simple way that has been suggested to resolve this issue is to use $2\alpha$ instead of $\alpha$ as the significance level, or equivalently, to divide the (two-sided) p-value obtained from the LR test by two [Pinheiro & Bates, 2000, Berkhof & Snijders, 2001, Cheung, 2014, 2015a]. That is, the null hypothesis is rejected at $\alpha=0.05$ if the calculated p-value is smaller than $0.102\alpha=$ 0.10, or equivalently, if the p-value divided by two is smaller than $\alpha=0.05$. In addition to testing the hypotheses, constructing confidence intervals for the variance components is of great interest. It has been suggested to use likelihood-based confidence intervals instead of Wald confidence intervals. In general, the likelihood-based approach for statistical inference about the variance parameters is preferred over Wald-type procedures, as it can account for the asymmetry or non-normality of the sampling distribution of the parameter estimates [Hox et al., 2010, Cheung, 2014, 2015a].
Having tested for and found evidence of heterogeneity among the true effect sizes, the logical next step in the random effects meta-analysis is to quantify such heterogeneity. In doing so, the $I^{2}$ index of traditional two-level meta-analysis, as given by eq. 10, can reasonably be extended to the three-level case. Thus, the $I^{2}$ measures at levels 2 and 3 can be calculated as
$I_{2}^{2}=\frac{\hat{\omega}^{2}}{S^{2}+\hat{\omega}^{2}+\hat{\tau}^{2}},$ (53)
and
$I_{3}^{2}=\frac{\hat{\tau}^{2}}{S^{2}+\hat{\omega}^{2}+\hat{\tau}^{2}},$ (54)
where $\hat{\omega}^{2}$ is an estimate of the amount of level 2 heterogeneity (i.e., within-study variance), $\hat{\tau}^{2}$ is an estimate of the amount of level 3 heterogeneity (i.e., between-study variance), and $S^{2}$ is an estimate of the typical within-study sampling variance ($\sigma^{2}$), as defined in eq. 11. Thus, $I_{2}^{2}$ and $I_{3}^{2}$ can be interpreted as the proportions of the total variation in observed effect sizes that are due to within-study heterogeneity (level 2) and between-study heterogeneity (level 3), respectively, rather than sampling error.
According to Cheung [2014, 2015a], the sampling variance $S^{2}$ depends on the sample sizes on which the effect size estimates are based and can therefore take different values for each set of studies selected. As $S^{2}$ is included in the calculation of $I_{2}^{2}$ and $I_{3}^{2}$, these indices depend on the sample sizes and thus do not estimate any population parameters. Consequently, Cheung [2014, 2015a] also define two intraclass correlations (ICCs) that include only the level-2 and level-3 variances. The ICC statistics are given as follows:
${ICC}_{2}^{2}=\frac{\hat{\omega}^{2}}{\hat{\omega}^{2}+\hat{\tau}^{2}}=\frac{I_{2}^{2}}{I_{2}^{2}+I_{3}^{2}},$ (55)
and
${ICC}_{3}^{2}=\frac{\hat{\tau}^{2}}{\hat{\omega}^{2}+\hat{\tau}^{2}}=\frac{I_{3}^{2}}{I_{2}^{2}+I_{3}^{2}},$ (56)
where ${ICC}_{2}^{2}$ and ${ICC}_{3}^{2}$ can be interpreted as the proportions of the total heterogeneity in the true effects attributed to level 2 (within-study heterogeneity) and level 3 (between-study heterogeneity), respectively. It is clear that these indices are independent of sample size.
The three-level random effects meta-analysis model can be extended to a mixed effects model by including effect size- and study-level characteristics as moderator variables ($x_{ij}$’s and $x_{j}$’s) in the Level 2 and Level 3 equations, respectively, to model the heterogeneity, if detected early. The resulting three-level equation can be expressed as follows:
$y_{ij}=\bm{x}_{ij}^{\prime}\bm{\beta}_{R}+u_{j}+\upsilon_{ij}+\varepsilon_{ij}.$ (57)
where $\bm{x}_{ij}$ is a $(p+1)\times 1$ vector of $p$ explanatory variables (which may include the study-level variables, $x_{j}$’s), plus a $1$ in the first position corresponding to the model intercept, and $\bm{\beta}_{R}$ is a $(p+1)\times 1$ vector of unknown constants to be estimated.
The model parameters (i.e., the variance components of the random effects and the fixed effects/regression coefficients) can be estimated simultaneously using maximum likelihood procedures (ML or REML). The variance components, denoted $\omega_{res}^{2}$ and $\tau_{res}^{2}$, represent the level 2 and level 3 residual (unexplained) heterogeneity, respectively. As in the two-level mixed effects model, the generalized Cochran’s $Q$ statistic given by eq. 20 (denoted $Q_{res}$) can be used to test the null hypothesis that the moderator variables included in the model fully explain the heterogeneity in the true effects, against the alternative that the moderators only partially explain the heterogeneity. It is also of interest to separately test whether the residual heterogeneity at each of the second and third levels is statistically different from zero. This can be done, as before, by using the likelihood ratio (LR) test for each hypothesis: $\text{H}_{0}:\omega_{res}^{2}=0$ and $\text{H}_{0}:\tau_{res}^{2}=0$ (as discussed earlier, this test is preferred over Wald-type tests). Rejection of the null hypothesis $\omega_{res}^{2}=0$ implies that the effect size-level covariates cannot fully explain the within-study heterogeneity in the true effects. Similarly, rejection of the null hypothesis $\tau_{res}^{2}=0$ implies that the study-level covariates cannot fully explain the between-study heterogeneity. Additionally, the $R^{2}$ discussed for two-level mixed effects meta-analysis can be extended to three-level meta-analysis, as follows:
$R_{2}^{2}=\frac{\hat{\omega}^{2}-\hat{\omega}_{res}^{2}}{\hat{\omega}^{2}}=1-\frac{\hat{\omega}_{res}^{2}}{\hat{\omega}^{2}},$ (58)
and
$R_{3}^{2}=\frac{\hat{\tau}^{2}-\hat{\tau}_{res}^{2}}{\hat{\tau}^{2}}=1-\frac{\hat{\tau}_{res}^{2}}{\hat{\tau}^{2}},$ (59)
where $\hat{\omega}^{2}$ and $\hat{\tau}^{2}$ denote the estimates of total within-study and between-study heterogeneity based on the meta-analysis model given by eq. 52, and $\hat{\omega}_{res}^{2}$ and $\hat{\tau}_{res}^{2}$ denote the estimates of within-study and between-study residual heterogeneity based on the meta-analysis model given by eq. 57. The $R_{2}^{2}$ index represents the proportion of the total within-study (level 2) heterogeneity that is explained by the effect size-level covariates included in the three-level mixed effects meta-analysis model, and the $R_{3}^{2}$ index represents the proportion of the total between-study (level 3) heterogeneity that is explained by the study-level covariates included in the model (see Cheung, 2014, 2015a).
It would also be of interest to test whether the coefficients on the moderator variables are statistically significant. A common procedure used in multilevel analyses is the LR test, which is conducted by comparing the log-likelihood values of the constrained (reduced) and unconstrained (full) models. This test can be performed with either ML or REML. As mentioned above, REML is generally preferred to ML as it provides more accurate estimates of variance components. However, this procedure only allows for comparison of models with the same fixed part (regression coefficients) and can thus be used to compare models that differ only in the random part (variance components). This means that the LR test for the significance of one or more of the slope coefficients can be performed with ML but not with REML. Similarly, the omnibus procedure is commonly used to test the null hypothesis that one or more of the slope coefficients are zero. Additionally, the Knapp-Hartung method can be used to test the statistical significance of a single slope coefficient, which is based on the t-distribution with an adjusted standard error.
The three-level meta-analysis can be conducted in R using the ‘$metafor$’ package [Viechtbauer, 2010, 2023]. Another option is the ‘$metaSEM$’ package [Cheung, 2015b], which offers a suite of functions for performing univariate, multivariate, and three-level meta-analyses through a structural equation modeling (SEM) approach [Cheung, 2013, 2014, 2015a]. [^11] Polanin et al. [2017] provide a review of 63 meta-analysis packages in R.
7 Concluding Remarks
Meta-analysis, a specific type of systematic review, is a robust statistical method used to combine quantitative findings from multiple studies, providing a comprehensive estimate of the overall or average effect size for a specific outcome of interest. The direction and magnitude of the estimated effect, accompanied by its confidence interval and hypothesis test results, provide insights into the phenomenon or relationship being investigated.
As a pivotal tool for synthesizing empirical evidence across various scientific disciplines, including economics, meta-analysis addresses the challenges posed by the exponential growth of research output. Individual studies often yield conflicting findings due to variations in data sources, methodologies, and contextual factors. By systematically integrating these results, meta-analysis bridges such gaps, offering a more robust and generalized understanding of economic phenomena. Its structured approach enhances statistical power, refines effect size estimates, and supports evidence-based decision-making.
A key strength of meta-analysis lies in its ability to address between-study heterogeneity—the differences in true effect sizes across studies. By employing meta-regression analysis, researchers can incorporate moderators representing study characteristics, thereby identifying the underlying sources of variability. This nuanced approach allows economists to draw more precise conclusions about the factors influencing economic relationships. For instance, it enables the examination of how varying labor market conditions shape the effectiveness of training programs or the differential impacts of policy interventions across regions and contexts.
[…]
Another critical aspect of meta-analysis is the treatment of effect size dependence. Effect size dependence refers to the statistical interdependence of effect sizes reported across or within studies, which violates the assumption of independence that underlies traditional meta-analytic methods. Failing to account for this dependence can lead to biased inferences. In economic datasets, multiple effect sizes are often reported within a single study, resulting in non-independence of observations. Advanced methods, such as multivariate meta-analysis and robust variance estimation, address this complexity by adjusting for the correlation between effect sizes, ensuring that the synthesized results provide a more accurate reflection of the true relationships under investigation.
This study reviews the key tools and techniques for applying meta-analysis in economics, with particular emphasis on addressing fundamental challenges such as publication bias, between-study heterogeneity, and effect size dependence. By exploring these methods, the study offers a valuable resource for researchers seeking to conduct robust meta-analyses in economics. It also demonstrates how advanced techniques, such as meta-regression, can yield actionable policy insights by uncovering the underlying factors contributing to variability in economic research. Ultimately, this work advances the methodological development of meta-analysis in economics, equipping researchers with the necessary knowledge to effectively apply meta-analytic methods, thereby enhancing the reliability of empirical findings and supporting evidence-based policy formulation.